### v2

# Algorithms for the Shortest Vector Problem in $2$-dimensional Lattices, Revisited

Abstract
Efficiently solving the Shortest Vector (Basis) Problem in two-dimensional lattices holds practical significance in cryptography and computational geometry. While simpler than its high-dimensional counterpart, two-dimensional SVP motivates scalable solutions for high-dimensional lattices and benefits applications like sequence cipher cryptanalysis involving large integers. In this work, we first propose a novel definition of reduced bases and develop an efficient adaptive lattice reduction algorithm CrossEuc that strategically applies the Euclidean algorithm across dimensions. Building on this framework, we introduce HVec, a vectorized generalization of the integer Half-GCD (HGCD) algorithm that efficiently halves vector bit-lengths, which may be of independent interest. By iteratively invoking HVec, our optimized algorithm HVecSBP achieves a reduced basis in $O(\log nM(n))$ time for arbitrary input bases with bit-length $n$, where $M(n)$ denotes the cost of multiplying two $n$-bit integers.
For finding the $\ell_{2}$-shortest basis, our new algorithms avoids costly inner products in its basic reduction step. The comprehensive experimental results demonstrate that (1) $\mathbf{CrossEuc}$ achieves approximately 500×500$\times$500 × and 350×350$\times$350 × speedup compared to Yap’s foundamental reduction algorithm $\mathbf{CRS}$ and the classical Lagrange reduction algorithm, respectively. (2) our optimized variant, $\mathbf{HVecSBP}$, delivers an additional 14$\times$ acceleration over Yap’s optimized $\mathbf{HalfGaussianSBP}$. For finding the $\ell_{\infty}$-shortest basis, our new designs eliminate Hermite Normal Form (HNF) conversion costs for general-form inputs and experimentally demonstrate significant improvements: (1) For HNF-form bases, $\mathbf{HVecSBP}$ achieves 13×13$\times$13 × acceleration over prior designs in worst-case scenarios; (2) For general-form bases,$\mathbf{CrossEuc}$ attains $\geq$4.5$\times$ efficiency gains over non-HGCD methods. Crucially, $\mathbf{HVecSBP}$ exhibits progressively intensifying advantages against HGCD-optimized designs as lattice basis vectors linear dependence weakens.
Index Terms:
Lattices, Half-GCD, Shortest Vector Problem (SVP), Hermite Normal Form (HNF), Lattice Basis Reduction
I Introduction
A lattice is a classic object of study in the geometry of numbers, originating from research on sphere packing and covering problems in the $17$th century. Around 1840, Gauss introduced the concept of a lattice and determined the maximum density of sphere packing in three-dimensional space. In the past thirty years, lattice theory has demonstrated its powerful applications in coding theory and cryptography, particularly through lattice-based cryptography. The application of lattice theory in these domains underscores its importance in tackling contemporary challenges in data security and communication. Its role in developing cryptographic methods that are resistant to both classical and quantum attacks highlights its value in securing digital information.
In general, a lattice is a set of discrete points in $n$-dimensional real space that has a periodic structure. Specifically, for $d$ linearly independent vectors $\bm{b}_{1},\cdots,\bm{b}_{d}$ in the $m$-dimensional real space $\mathbb{R}^{m}$, a lattice $\mathcal{L}$
$\mathcal{L}=L(\bm{B})=\left\{\sum_{i=1}^{d}z_{i}\bm{b}_{i}:z_{i}\in\mathbb{Z}\right\}.$
is the set of all integer linear combinations of these vectors, where $d$ is called the dimension of the lattice and the set of $d$ vectors $\bm{B}=[\bm{b}_{1},\cdots,\bm{b}_{d}]$ is called a basis of the lattice. The $k$-th successive minimum of a lattice is defined as
$\lambda_{k}(\mathcal{L})=\mathrm{inf}\{r>0:\mathrm{dim}(\mathrm{span}(\mathcal{B}(\bm{o},r)\cap\mathcal{L}))\geq k\},$
which $\mathcal{B}(\bm{o},r)$ represents the open ball of radius $r$ centered at the origin. In particular, $\lambda_{1}(L)$ denotes the length of the shortest non-zero vector in the lattice.
One of the most fundamental computational problems in lattice theory, known as the Shortest Vector Problem (SVP), is to find the shortest vector in a lattice. For a given lattice $\mathcal{L}$, this involves identifying a non-zero lattice vector $\mathbf{v}$ such that $||\mathbf{v}||\leq||\mathbf{u}||$ holds for all non-zero vectors $\mathbf{u}\in\mathcal{L}$. The difficulty of solving this problem grows significantly with the dimension of the lattice. However, in the simplest case of two-dimensional lattices, solving the SVP—or even the Shortest Basis Problem (SBP), which seeks a basis $\mathbf{B}=[\mathbf{a}\ \mathbf{b}]$ satisfying $\|\mathbf{a}\|=\lambda_{1}(\mathcal{L})$ and $\|\mathbf{b}\|=\lambda_{2}(\mathcal{L})$-is relatively straightforward, making it an ideal starting point for studying lattice problems. Developing fast algorithms for solving the two-dimensional SVP/SBP is of both theoretical and practical importance. From an algorithm design perspective, solving the SVP in two-dimensional lattices provides foundational insights for addressing higher-dimensional cases. For example, Lagrange’s pioneering work [6] on two-dimensional lattice reduction laid the groundwork for the celebrated LLL algorithm by A. K. Lenstra, H. W. Lenstra, and L. Lovász [7]. The LLL algorithm extends Lagrange’s approach by introducing techniques for reducing lattice bases in higher dimensions, combining ideas of approximate orthogonality and length minimization. As such, understanding and refining two-dimensional lattice reduction algorithms can directly contribute to advancements in higher-dimensional lattice computations. On the practical side, the SVP in two-dimensional lattices has direct applications in cryptoanalysis. One critical indicator of the security of cryptographic random sequences is their $2$-adic complexity, which measures the sequence’s resistance to linear feedback shift register attacks. Computing the $2$-adic complexity involves determining the Minimal Rational Fractional Representation (MRFR) of the sequence, a problem that can be directly transformed into finding the shortest basis of a two-dimensional lattice in the $\ell_{\infty}$ norm. Specifically, given a sequence, a specialized two-dimensional lattice is constructed in Hermite Normal Form (HNF), and the shortest vector in this lattice corresponds to the MRFR. Efforts to optimize algorithms for solving the MRFR problem have shown significant progress. Arnault et al. [2] proposed a Euclidean-based algorithm in 2004, with subsequent improvements in 2008 [1], to compute the shortest basis of the lattice. However, these methods occasionally produced incorrect results in certain cases. Recently, Che et al. [3] proposed a novel and faster algorithm targeting the two-dimensional lattice shortest vector problem, significantly enhancing both the computational efficiency and reliability of solutions. Further optimization of algorithms for finding the shortest basis in two-dimensional lattices would directly accelerate the computation of $2$-adic complexity, enabling faster and more accurate assessments of cryptographic random sequence security.
In summary, the above discussion highlights the two-fold significance of studying two-dimensional lattice problems: as a foundation for high-dimensional lattice theory and as a means to improve practical cryptographic applications. Therefore, continued research in this area is vital for advancing both theoretical lattice studies and practical algorithmic solutions.
I-A Related works
[…]
to minimize the length of $\bm{c}=\bm{a}-q\bm{b}$, updating the basis to $[\bm{b},\bm{c}]$. Iterating this process yields the shortest basis in $O(M(n)n)$ time, where $n=\max\{\lceil\log\|\bm{a}\|\rceil,\lceil\log\|\bm{b}\|\rceil\}$ and $M(n)$ is the complexity of multiplying two $n$-bit integers.
For the frequently used Euclidean metric ($\ell_{2}$ norm), Lagrange reduction uses:
$q=\left\lceil\frac{\langle\bm{a},\bm{b}\rangle}{\langle\bm{b},\bm{b}\rangle}\right\rfloor.$
This $q$-selection differs from the standard integer Euclidean algorithm. Yap [15] studied a directly analogous approach, updating the basis to $[\bm{b},\bm{c}]$ with $\bm{c}=\bm{a}-q\bm{b}$ via:
$q=\left\lfloor\frac{\langle\bm{a},\bm{b}\rangle}{\langle\bm{b},\bm{b}\rangle}\right\rfloor.$
The algorithm terminates if the basis remains unchanged after an update. This method, termed $\mathbf{CRS}$, was proven to have slightly higher complexity than Lagrange’s reduction, but with negligible efficiency loss (”not giving up too much” as said in [15]). Crucially, it preserves the geometric property that basis vectors always form acute angles. Further optimizing $\mathbf{CRS}$, Yap designed a Half-Gaussian algorithm for integer vectors inspired by the integer HGCD method, reducing the time complexity to $O(M(n)\log n)$. Rote’s algorithm [13] leverages Yap’s lattice basis reduction [15] to solve SVP for a two-dimensional lattice module $m$ in $O(\log m(\log\log m)^{2})$ time.
For the special Minkowski metric ($\ell_{\infty}$ norm), Eisenbrand [4] and Che et al. [3] subsequently investigated fast algorithms for two-dimensional lattices with a known HNF basis. Specifically, when the input lattice basis satisfies $a_{1}>b_{1}\geq 0$, $a_{2}=0$ and $b_{2}\neq 0$. it was noted that by taking
$q=\left\lfloor\frac{a_{1}}{b_{1}}\right\rfloor$
the shortest vector can be found more efficiently than with the Lagrange reduction algorithm. In contrast to Eisenbrand’s method [4] which addresses the standard HNF, Che et al. [3] specifically target a specialized HNF variant characterized by the constraint $b_{2}=1$. Notably, Che et al. [3] formalize an exact termination criterion for basis reduction and propose an optimized algorithm through systematic integration of the integer HGCD algorithm. Their implementation achieves a time complexity of $O(M(n)\log n)$, supported by comprehensive technical details. In comparison, Eisenbrand [4] interprets their algorithm through continued fraction theory and theoretically suggests potential optimizations via the integer HGCD algorithm to attain the same asymptotic complexity $O(M(n)\log n)$, though without providing explicit implementation specifics. Wu and Xu [14] linked Qin Jiushao’s algorithm with the solution of SVP in a special kind of two-dimensional lattices with $a_{1}=1$, $b_{1}=0$ and $\gcd(a_{2},b_{2})=1$, and proposed an efficient method to solve the SVP of this kind of two-dimensional lattices with
$q=\left\lfloor\frac{a_{1}-1}{b_{1}}\right\rfloor.$
By examining each state matrix generated during the execution of the algorithm, the two vectors in the current state and their simple combinations can be utilized to extract the shortest lattice vector.
Regarding the relationship between the shortest vector under the $\ell_{\infty}$ norm and that under the $\ell_{2}$ norm, Eisenbrand [4] cites Lagarias’ work [5] to demonstrate that reduced lattice basis containing the $\ell_{\infty}$-shortest vector can be reduced to the $\ell_{2}$-shortest vector through a constant number of Lagrange reductions.
I-B Motivation and Our Contribution
Existing research indicates that the SVP (SBP) in two-dimensional lattices remains imperfectly resolved. Under the commonly used Euclidean metric, Although Yap’s Half-Gaussian algorithm [15] achieves lower complexity than Lagrange reduction, its underlying fundamental reduction algorithm (CRS) – as Yap himself acknowledges – is not inherently more efficient than classical Lagrange reduction. This naturally raises a pivotal question: Does there exist a fundamental reduction algorithm that is more efficient than the Lagrange reduction algorithm, along with an optimized version that can outperform the Half-Gaussian algorithm in finding the $\ell_{2}$-shortest vector (basis)? For the special Minkowski metric, current algorithms face the following limitations:
  • For HNF-form bases $\mathbf{B}=\begin{pmatrix}a_{1}&b_{1}\\ 0&b_{2}\end{pmatrix}$: While prior works suggest that integer HGCD algorithms can optimize continued fraction-based methods, detailed technical analyses and implementations remain scarce. As noted in [11], “the integer HGCD algorithm is intricate, error-prone, and rarely fully detailed”, and [12] emphasizes its implementation challenges due to “numerous sub-cases and limited practical adoption.” Notably, [3] provides a concrete HGCD-based SVP solution for the special case $b_{2}=1$, but the implementation details for general $b_{2}$ is not available.
  • For general-form bases: Current algorithms first convert inputs to HNF via (HGCD-based) extended Euclidean algorithms. However, this conversion often inflates integer sizes, especially for short initial bases, leading to longer lattice vectors. This raises two interesting questions: (i) Is HNF conversion necessary? (ii) Can direct reduction algorithms bypass this step for a faster performance?
To address these challenges, this work aims to develop a novel fundamental reduction algorithm achieving: (1) For Euclidean metric: Superior efficiency over both Lagrange reduction and Yap’s CRS, with an optimized version surpassing Half-Gaussian. (2) For Minkowski metric: Elimination of HNF conversion and generalization of HGCD-based optimization to arbitrary $b_{2}$, thereby bridging the gap between theoretical possibilities and practical implementations. Concretely, our main contributions can be summarized as follows:
  1. 1.
    Novel Definition of a Reduced Basis. For the first time, we explicitly propose a novel definition of a reduced basis distinct from the conventional Lagrange-reduced basis. This newly defined basis not only contains the $\ell_{\infty}$-shortest vector but also yields the $\ell_{2}$-shortest vector in at most one additional addition/subtraction step. This framework extends the theoretical foundation of lattice reduction and is expected to provide a solid foundation for further research in this field.
  2. 2.
    Efficient Algorithms with Detailed Implementation and Rigorous Complexity Analysis. For our newly introduced reduced basis, we design a novel fundamental algorithm named CrossEuc that bypasses HNF conversion, achieving a complexity of $O(n^{2})$. This approach eliminates the computational overhead associated with HNF transformation. Furthermore, we introduce a vectorized adaptation of the HGCD algorithm, termed HVec, which represents the first explicit extension of HGCD-like techniques to vector pairs. This innovation is of independent theoretical interest and significantly advances the efficiency of solving the SVP in two-dimensional lattices. Building on this, we present an optimized version HVecSBP of the fundamental algorithm CrossEuc that reduces the complexity to $O(\log n\cdot M(n))$, where $M(n)$ denotes the cost of multiplying two $n$-bit integers.
  3. 3.
    Comprehensive Experimental Performance Analysis. We conduct extensive experiments comparing our methods against state-of-the-art algorithms. Results show that (1) For computing the $\ell_{2}$-shortest basis with input lattice basis size $2\times 10^{5}$, regardless of whether the input is in HNF form or general form, our proposed fundamental reduction algorithm, $\mathbf{CrossEuc}$, achieves approximately 500×500$\times$500 × and 350×350$\times$350 × speedup compared to Yap’s foundamental reduction algorithm $\mathbf{CRS}$ and the classical Lagrange reduction algorithm, respectively. In comparison with Yap’s optimized algorithm, $\mathbf{HalfGaussianSBP}$, our optimized variant, $\mathbf{HVecSBP}$, achieves an additional speedup of approximately 14×14$\times$14 ×. (2) For computing the $\ell_{\infty}$-shortest basis with input lattice basis size $10^{6}$, our new algorithms demonstrate significant improvements: (2.1) For HNF-form bases, $\mathbf{HVecSBP}$ achieves 13×13$\times$13 × acceleration over prior designs in worst-case scenarios. (2.2) For general-form bases, compared to the previous designs without using HGCD optimization, the proposed algorithm CrossEuc achieves a at least 4.5×4.5$\times$4.5 × efficiency improvement. Compared to the previous designs using HGCD optimization, when the input vectors are nearly degenerate (i.e., treated as integers), HGCD-HNF-HVecSBP method retain a slight advantage. As the linear dependency between vectors weakens, HVecSBP exhibits remarkably growing efficiency gains.
I-C Road Map
The paper is organized as follows. Section II introduces the notations, terminologies, and the mathematical concepts and properties frequently referenced in this work. In Section III, we define our proposed reduced basis and establish its theoretical properties. Section IV presents the CrossEuc algorithm alongside a detailed analysis. In Section V, we propose the HVec algorithm, discuss its analysis, and then introduce the HVecSBP algorithm, which iteratively invokes HVec to solve the shortest basis problem in two-dimensional lattices. Section VI provides a comprehensive experimental evaluation comparing the practical performance of our improved algorithms with that of existing methods, and the paper concludes with a brief summary in the final section.
II Preliminaries
For completeness, this section introduces the necessary preliminaries for the design and analysis of our new algorithm.
II-A Notations and Terminologies
Throughout this paper, $\mathbb{Z}$ denotes the ring of integers, and $\mathbb{R}$ represents the field of real numbers. For $x,y\in\mathbb{R}$, #$\#x=\left\lceil\log_{2}(|x|+1)\right\rceil$ denotes the bit size of $x$, while #$\#(x,y)=\max\{\#x,\#y\}$ and #¯$\underline{\#}(x,y)=\min\{\#x,\#y\}$. The round function $\left\lceil x\right\rfloor$ represents the nearest integer to $x$, the floor function $\left\lfloor x\right\rfloor$ represents the greatest integer less than or equal to $x$, and the ceiling function $\left\lceil x\right\rceil$ represents the smallest integer greater than or equal to $x$. The ”rounding towards zero” $\left\lfloor x\right\rfloor_{o}$ of $x$ and the sign function $\mathsf{sgn}(x)$ of $x$ are defined as follows:
$\left\lfloor x\right\rfloor_{o}=\begin{cases}\left\lfloor x\right\rfloor,&\text{if }x\geq 0\\ \left\lceil x\right\rceil,&\text{else}\end{cases},\ \mathsf{sgn}(x)=\begin{cases}1,&\text{if }x\geq 0\\ -1,&\text{else}\end{cases}.$
We use $M(n)$ to denote the time complexity of the multiplication of two $n$-bit integers. Uppercase bold letters represent matrices, while lowercase bold letters represent vectors. For any two vectors $\mathbf{a}=(a_{1},a_{2})^{T}$ and $\mathbf{b}=(b_{1},b_{2})^{T}$, we define #$\#\mathbf{a}=\max\{\#a_{1},\#a_{2}\}$, #$\#(\mathbf{a},\mathbf{b})=\max\{\#\mathbf{a},\#\mathbf{b}\}$ and #¯$\underline{\#}(\mathbf{a},\mathbf{b})=\min\{\#\mathbf{a},\#\mathbf{b}\}$. For a vector $\bm{x}=(x_{1},x_{2},\ldots,x_{m})\in\mathbb{R}^{m}$, its Euclidean length ($\ell_{2}$ norm) is defined as $\|\bm{x}\|=\sqrt{\sum_{i=1}^{m}x_{i}^{2}}$, and its Minkowski length ($\ell_{\infty}$ norm) is defined as $\|\bm{x}\|_{\infty}=\max_{1\leq i\leq m}|x_{i}|$.
II-B Lattices and related properties
In this section, we introduce the definition of lattices and some necessary properties we used in the rest of the paper.
Definition 1 ([9]).
(Lattice). Given linearly independent vectors $\bm{b_{1}},\bm{b_{2}},...,\bm{b_{d}}\in\mathbb{R}^{m}$, the lattice generated by these vectors is defined as
$\mathcal{L}=L(\bm{B})=L(\bm{b_{1}},\bm{b_{2}},...,\bm{b_{d}})=\left\{\sum_{i=1}^{d}z_{i}\bm{b}_{i}:z_{i}\in\mathbb{Z}\right\},$
where $d$ and $m$ are called the rank and the dimension of the lattice, respectively. Specially, if $d=m$, the lattice is called full-rank.
Definition 2 ([9]).
(Determinant). Let $\mathcal{L}=L(\mathbf{B})$ be a lattice of rank $d$. The determinant of $\mathcal{L}$, denoted $\det(\mathcal{L})=\sqrt{\det\left(\mathbf{B}^{T}\mathbf{B}\right)}$. Specially, if $\mathcal{L}$ is full-rank, $\det(\mathcal{L})=|\det(\mathbf{B})|$.
Definition 3 ([9]).
(Unimodular Matrix). A matrix $\mathbf{U}\in\mathbb{Z}^{d\times d}$ is called unimodular if $\det(\mathbf{U})=\pm 1$.
Lemma 1 ([9]).
Two bases $\mathbf{B_{1}},\mathbf{B_{2}}\in\mathbb{R}^{m\times d}$ are equivalent if and only if $\mathbf{B_{2}}=\mathbf{B_{1}}\mathbf{U}$ for some unimodular matrix $\mathbf{U}$.
Definition 4 ([9]).
(Successive Minima). Let $\mathcal{L}$ be a lattice of rank $d$. For $k\in\left\{1,...,d\right\}$, the $k$th successive minimum is defined as
$\lambda_{k}(\mathcal{L})=\mathrm{inf}\{r>0:\mathrm{dim}(\mathrm{span}(\mathcal{B}(\bm{o},r)\cap\mathcal{L}))\geq k\},$
[…]
For two-dimensional lattices, Lagrange [6] introduced the concept of a reduced basis for two-dimensional lattices. That is,
Definition 5 ([6, 9]).
[…]
For a Lagrange-reduced basis, the following result is well-known:
Lemma 2 ([9]).
[…]
The above property is derived from the following general fact:
Lemma 3 ([9]).
[…]
II-C Hermite Normal Form
The Hermite Normal Form (HNF) of a matrix is a canonical form used in linear algebra and number theory for integer matrices.
Definition 6 ([8]).
(Hermite Normal Form). A nonsingular matrix $\mathbf{H}\in\mathbb{Z}^{d\times d}$ is said to be in HNF if (1) $h_{i,i}>0$ for $1\leq i\leq d$. (2) $h_{j,i}=0$ for $i<j\leq d$. (3) $0\leq h_{j,i}<h_{i,i}$ for $1\leq j<i$.
For any nonsingular integer matrix $\mathbf{A}$, there exists a unique matrix $\mathbf{H}$ in Hermite Normal Form (HNF) and a unique unimodular matrix $\mathbf{U}$ such that $\mathbf{H}=\mathbf{U}\mathbf{A}$ [8]. Specifically, for any two-dimensional lattice, the basis matrix can be efficiently converted into its HNF using the following lemma.
Lemma 4 ([4]).
Given a matrix $\mathbf{B}=\begin{pmatrix}a_{1}&b_{1}\\ a_{2}&b_{2}\end{pmatrix}\in\mathbb{Z}^{2\times 2}$, let $c=\gcd(a_{2},b_{2})=xa_{2}+yb_{2}$ be the greatest common divisor of $a_{2}$ and $b_{2}$, then
$\begin{pmatrix}a_{1}&b_{1}\\ a_{2}&b_{2}\end{pmatrix}\begin{pmatrix}b_{2}/\gcd(a_{2},b_{2})&x\\ -a_{2}/\gcd(a_{2},b_{2})&y\end{pmatrix}=\begin{pmatrix}a&b\\ 0&c\end{pmatrix}\in\mathbb{Z}^{2\times 2}$
with $a=(a_{1}b_{2}-a_{2}b_{1})/\gcd(a_{2},b_{2})$ and $b=a_{1}x+b_{1}y$. By applying elementary vector transformations, we can ensure that the matrix satisfy the HNF conditions.
III Our new reduced basis and its properties
In this section, we first introduce a new reduced basis and then present its properties. Throughout this section and the rest of the paper, unless otherwise specified, the notations $\|\mathbf{x}\|$ and $\|\mathbf{x}\|_{2}$ are used to represent the Minkowski metric ($\ell_{\infty}$ norm) of $\mathbf{x}$ and the Euclidean metric ($\ell_{2}$ norm), respectively.
Definition 7.
Given a lattice $\mathcal{L}=L(\bm{B})$ with a basis
$\mathbf{B}=[\mathbf{a}\ \mathbf{b}]=\begin{pmatrix}a_{1}&b_{1}\\ a_{2}&b_{2}\end{pmatrix}\in\mathbb{R}^{2\times 2},$ (1)
we call $\mathbf{B}$ is reduced if
$a_{1}a_{2}b_{1}b_{2}\leq 0\wedge(|a_{1}|-|a_{2}|)(|b_{1}|-|b_{2}|)\leq 0$ (2)
It should be remarked that the equation (2) also can be equivalently reformulated as $a_{1}a_{2}b_{1}b_{2}\leq 0\wedge|a_{1}b_{1}-a_{2}b_{2}|\leq|\det(\mathbf{B})|.$ Now we present two important properties of the new defined reduced basis. The first theorem establishes the minimality of the reduced basis under the Minkowski ($\ell_{\infty}$) norm, while the second theorem demonstrates a similar property under the Euclidean ($\ell_{2}$) norm.
Theorem 1.
Let $\mathcal{L}=L(\bm{B})$ be a lattice with basis $\mathbf{B}$ defined in equation (1), and let $\lambda_{1}(\mathcal{L})$ and $\lambda_{2}(\mathcal{L})$ denote the successive minima under $\ell_{\infty}$ norm. If $\mathbf{B}$ is reduced, then $\lambda_{1}(\mathcal{L})=\min\{\|\mathbf{a}\|,\|\mathbf{b}\|\}.$ Further, let $\mathbf{c}=(c_{1}\ c_{2})\in\mathcal{L}$ be a vector achieving $\lambda_{2}(\mathcal{L})$. Then $\mathbf{c}=\mathrm{argmax}\{\|\mathbf{a}\|,\|\mathbf{b}\|\}-z\cdot\mathrm{argmin}\{\|\mathbf{a}\|,\|\mathbf{b}\|\}$, where $z$ is determined as follows:
(1) If $a_{1}b_{1}\geq 0,a_{2}b_{2}\leq 0$, then
$z=\left\{\begin{array}[]{ll}\left\lfloor\frac{|b_{1}|-|b_{2}|}{|a_{1}|+|a_{2}|}\right\rfloor\ \mbox{or}\ \left\lceil\frac{|b_{1}|-|b_{2}|}{|a_{1}|+|a_{2}|}\right\rceil,&\|\mathbf{a}\|\leq\|\mathbf{b}\|\\ \left\lfloor\frac{|a_{1}|-|a_{2}|}{|b_{1}|+|b_{2}|}\right\rfloor\ \mbox{or}\ \left\lceil\frac{|a_{1}|-|a_{2}|}{|b_{1}|+|b_{2}|}\right\rceil,&\|\mathbf{a}\|>\|\mathbf{b}\|\end{array}\right.$
(2) If $a_{1}b_{1}<0,a_{2}b_{2}\geq 0$, then
$z=\left\{\begin{array}[]{ll}\left\lfloor\frac{|b_{2}|-|b_{1}|}{|a_{1}|+|a_{2}|}\right\rfloor\ \mbox{or}\ \left\lceil\frac{|b_{2}|-|b_{1}|}{|a_{1}|+|a_{2}|}\right\rceil,&\|\mathbf{a}\|\leq\|\mathbf{b}\|\\ \left\lfloor\frac{|a_{2}|-|a_{1}|}{|b_{1}|+|b_{2}|}\right\rfloor\ \mbox{or}\ \left\lceil\frac{|a_{2}|-|a_{1}|}{|b_{1}|+|b_{2}|}\right\rceil,&\|\mathbf{a}\|>\|\mathbf{b}\|\end{array}\right.$
Specially, if $a_{1}a_{2}b_{1}b_{2}=0$ and none of the above cases apply, negate $\mathbf{a}$ or $\mathbf{b}$ as necessary to fit into one of the above cases.
Proof.
Without loss of generality, we only consider the case $a_{1}b_{1}\geq 0$ and $a_{2}b_{2}\leq 0$. We aim to prove that for any non-zero lattice vector $\mathbf{v}=z_{1}\mathbf{a}+z_{2}\mathbf{b}=(z_{1}a_{1}+z_{2}b_{1}\ z_{1}a_{2}+z_{2}b_{2})^{T}$, the following inequality holds:
$\min\{\|\mathbf{a}\|,\|\mathbf{b}\|\}\leq\|\mathbf{v}\|.$ (3)
Clearly, in case that $z_{1}=0,z_{2}\neq 0$,
$\|\mathbf{v}\|=\max\{|z_{1}a_{1}+z_{2}b_{1}|,|z_{1}a_{2}+z_{2}b_{2}|\}=\max\{|z_{2}b_{1}|,|z_{2}b_{2}|\}\geq\max\{|b_{1}|,|b_{2}|\}=\|\mathbf{b}\|\geq\min\{\|\mathbf{a}\|,\|\mathbf{b}\|\}.$
and, in case that $z_{1}\neq 0,z_{2}=0$,
$\|\mathbf{v}\|=\max\{|z_{1}a_{1}+z_{2}b_{1}|,|z_{1}a_{2}+z_{2}b_{2}|\}=\max\{|z_{1}a_{1}|,|z_{1}a_{2}|\}\geq\max\{|a_{1}|,|a_{2}|\}=\|\mathbf{a}\|\geq\min\{\|\mathbf{a}\|,\|\mathbf{b}\|\}.$
We now analyze the case $z_{1}z_{2}\neq 0$ based on the properties of the reduced basis ($i.e.,(|a_{1}|-|a_{2}|)(|b_{1}|-|b_{2}|)\leq 0$).
  1. 1.
    $|a_{1}|=|a_{2}|$. Here, $\|\mathbf{a}\|=\max\{|a_{1}|,|a_{2}|\}=|a_{1}|=|a_{2}|$. Then, if $z_{1}z_{2}>0$, we have $\|\mathbf{v}\|=\max\{|z_{1}a_{1}+z_{2}b_{1}|,|z_{1}a_{2}+z_{2}b_{2}|\}\geq|z_{1}a_{1}+z_{2}b_{1}|\geq\max\{|a_{1}|,|b_{1}|\}\geq|a_{1}|=\|\mathbf{a}\|\geq\min\{\|\mathbf{a}\|,\|\mathbf{b}\|\}$ and, if $z_{1}z_{2}<0$, we have $\|\mathbf{v}\|=\max\{|z_{1}a_{1}+z_{2}b_{1}|,|z_{1}a_{2}+z_{2}b_{2}|\}\geq|z_{1}a_{2}+z_{2}b_{2}|\geq|a_{2}|=\|\mathbf{a}\|\geq\min\{\|\mathbf{a}\|,\|\mathbf{b}\|\}.$
  2. 2.
    $|b_{1}|=|b_{2}|$. Here, $\|\mathbf{b}\|=\max\{|b_{1}|,|b_{2}|\}=|b_{1}|=|b_{2}|$. Then, if $z_{1}z_{2}>0$, we have $\|\mathbf{v}\|=\max\{|z_{1}a_{1}+z_{2}b_{1}|,|z_{1}a_{2}+z_{2}b_{2}|\}\geq|z_{1}a_{1}+z_{2}b_{1}|\geq\max\{|a_{1}|,|b_{1}|\}\geq|b_{1}|=\|\mathbf{b}\|\geq\min\{\|\mathbf{a}\|,\|\mathbf{b}\|\}$ and, if $z_{1}z_{2}<0$, we have $\|\mathbf{v}\|=\max\{|z_{1}a_{1}+z_{2}b_{1}|,|z_{1}a_{2}+z_{2}b_{2}|\}\geq|z_{1}a_{2}+z_{2}b_{2}|\geq|b_{2}|=\|\mathbf{b}\|\geq\min\{\|\mathbf{a}\|,\|\mathbf{b}\|\}.$
  3. 3.
[…]
  4. 4.
    $|a_{1}|>|a_{2}|$ and $|b_{1}|<|b_{2}|$. Here, $\|\mathbf{a}\|=\max\{|a_{1}|,|a_{2}|\}=|a_{1}|$ and $\|\mathbf{b}\|=\max\{|b_{1}|,|b_{2}|\}=|b_{2}|$. Then, if $z_{1}z_{2}>0$, we have $\|\mathbf{v}\|=\max\{|z_{1}a_{1}+z_{2}b_{1}|,|z_{1}a_{2}+z_{2}b_{2}|\}\geq|z_{1}a_{1}+z_{2}b_{1}|\geq\max\{|a_{1}|,|b_{1}|\}\geq|a_{1}|=\|\mathbf{a}\|\geq\min\{\|\mathbf{a}\|,\|\mathbf{b}\|\}$,and, if $z_{1}z_{2}<0$, we have $\|\mathbf{v}\|=\max\{|z_{1}a_{1}+z_{2}b_{1}|,|z_{1}a_{2}+z_{2}b_{2}|\}\geq|z_{1}a_{2}+z_{2}b_{2}|\geq|b_{2}|=\|\mathbf{b}\|\geq\|\mathbf{a}\|\geq\min\{\|\mathbf{a}\|,\|\mathbf{b}\|\}.$
Overall, the equation (3) holds, confirming that $\lambda_{1}(\mathcal{L})=\min\{\|\mathbf{a}\|,\|\mathbf{b}\|\}$. Next, we will analyze the vector that achieves the second successive minimum $\lambda_{2}(\mathcal{L})$, by considering two distinct cases.
Case 1: $\min\{\|\mathbf{a}\|,\|\mathbf{b}\|\}=\|\mathbf{a}\|=\lambda_{1}(\mathcal{L})$, i.e., $\max\{|a_{1}|,|a_{2}|\}\leq\max\{|b_{1}|,|b_{2}|\}$. Define the function
$f(x)=\|\mathbf{b}-x\mathbf{a}\|=\max\{|b_{1}-xa_{1}|,|b_{2}-xa_{2}|\}\triangleq\max\{f_{1}(x),f_{2}(x)\}.$
Let $f(z)=\min_{z\in\mathbb{Z}}f(x)$. Then, $\|\mathbf{a}\|\leq f(z)=\|\mathbf{b}-z\mathbf{a}\|\leq f(z+1)=\|\mathbf{b}-(z+1)\mathbf{a}\|=\|\mathbf{a}-(\mathbf{b}-z\mathbf{a})\|$ and $f(z-1)=\|\mathbf{b}-(z-1)\mathbf{a}\|=\|\mathbf{a}+(\mathbf{b}-z\mathbf{a})\|$. By Lemma 2, it follows that $\lambda_{2}(\mathcal{L})=f(z)$. Thus, we only need to estimate $z$. We now consider the case where $a_{1}a_{2}\neq 0$. The argument is essentially the same for cases where $a_{1}=0$ or $a_{2}=0$. From the definitions of $f_{1}(x)$ and $f_{2}(x)$, we have
$f_{1}(x)=\left\{\begin{array}[]{ll}|a_{1}|x-|b_{1}|,&x\geq\frac{b_{1}}{a_{1}}\\ -|a_{1}|x+|b_{1}|,&x<\frac{b_{1}}{a_{1}}\end{array}\right.,\ f_{2}(x)=\left\{\begin{array}[]{ll}|a_{2}|x+|b_{2}|,&x\geq\frac{b_{2}}{a_{2}}\\ -|a_{2}|x-|b_{2}|,&x<\frac{b_{2}}{a_{2}}\end{array}\right..$
The minimum $f(z)$ is achieved at the integer closest to the point of intersection where $-|a_{1}|x+|b_{1}|=|a_{2}|x+|b_{2}|$ (i.e., $x=\frac{|b_{1}|-|b_{2}|}{|a_{1}|+|a_{2}|}$). Therefore,
$z=\left\lfloor\frac{|b_{1}|-|b_{2}|}{|a_{1}|+|a_{2}|}\right\rfloor\ \mbox{or}\ \left\lceil\frac{|b_{1}|-|b_{2}|}{|a_{1}|+|a_{2}|}\right\rceil.$
Case 2: $\min\{\|\mathbf{a}\|,\|\mathbf{b}\|\}=\|\mathbf{b}\|$, i.e., $\max\{|a_{1}|,|a_{2}|\}\geq\max\{|b_{1}|,|b_{2}|\}$. Define the function
$g(y)=\|\mathbf{a}-y\mathbf{b}\|=\max\{|a_{1}-yb_{1}|,|a_{2}-yb_{2}|\}\triangleq\max\{g_{1}(y),g_{2}(y)\}.$
Let $g(z)=\min_{z\in\mathbb{Z}}g(y)$. Then $\|\mathbf{b}\|\leq g(z)=\|\mathbf{a}-z\mathbf{b}\|\leq g(z+1)=\|\mathbf{a}-(z+1)\mathbf{b}\|=\|\mathbf{b}-(\mathbf{a}-z\mathbf{b})\|$ and $g(z-1)=\|\mathbf{b}-(z-1)\mathbf{a}\|=\|\mathbf{b}+(\mathbf{a}-z\mathbf{b})\|$. By Lemma 2, it follows that $\lambda_{2}(\mathcal{L})=g(z)$. Thus, we only need to estimate $z$. We now consider the case where $b_{1}b_{2}\neq 0$. The argument is essentially the same for cases where $b_{1}=0$ or $b_{2}=0$. From the definitions of $g_{1}(y)$ and $g_{2}(y)$, we have
$g_{1}(y)=\left\{\begin{array}[]{ll}|b_{1}|y-|a_{1}|,&y\geq\frac{a_{1}}{b_{1}}\\ -|b_{1}|y+|a_{1}|,&y<\frac{a_{1}}{b_{1}}\end{array}\right.,\ g_{2}(y)=\left\{\begin{array}[]{ll}|b_{2}|y+|a_{2}|,&y\geq\frac{a_{2}}{b_{2}}\\ -|b_{2}|y-|a_{2}|,&y<\frac{a_{2}}{b_{2}}\end{array}\right..$
The minimum $g(z)$ will achieve at the integer around the point of intersection $-|b_{1}|y+|a_{1}|=|b_{2}|y+|a_{2}|$ (i.e., $y=\frac{|a_{1}|-|a_{2}|}{|b_{1}|+|b_{2}|}$). Therefore,
$z=\left\lfloor\frac{|a_{1}|-|a_{2}|}{|b_{1}|+|b_{2}|}\right\rfloor\ \mbox{or}\ \left\lceil\frac{|a_{1}|-|a_{2}|}{|b_{1}|+|b_{2}|}\right\rceil.$
Theorem 2.
Let $\mathcal{L}=L(\bm{B})$ be a lattice with basis $\mathbf{B}$ defined in equation (1), and let $\lambda_{1}(\mathcal{L})$ and $\lambda_{2}(\mathcal{L})$ denote the successive minims under $\ell_{2}$ norm. If $\mathbf{B}$ is reduced, then $\mathbf{u}=\mathrm{argmin}\{\|\mathbf{a}\|_{2},\|\mathbf{b}\|_{2},\|\mathbf{a}+\mathbf{b}\|_{2},\|\mathbf{a}-\mathbf{b}\|_{2}\}$ is a shortest non-zero lattice vector under the $\ell_{2}$ norm. That is, $\lambda_{1}(\mathcal{L})=\|\mathbf{u}\|_{2}$. Further, for any vector $\mathbf{x}\in\{\mathbf{a},\mathbf{b}\}\setminus\{\mathbf{u}\}$, $\mathbf{v}=\mathbf{x}-q\mathbf{u}$ with $q=\left\lceil\frac{\langle\mathbf{x},\mathbf{u}\rangle}{\langle\mathbf{u},\mathbf{u}\rangle}\right\rfloor$ is a lattice vector achieving $\lambda_{2}(\mathcal{L})$. That is, $\lambda_{2}(\mathcal{L})=\|\mathbf{v}\|_{2}$.
Proof.
First, we prove $\lambda_{1}(\mathcal{L})=\|\mathbf{u}\|_{2}$. Namely, we need to prove $\forall(z_{1}\ z_{2})\in\mathbb{Z}^{2}\backslash\{\mathbf{0}\},\|\mathbf{v}\|_{2}=\|z_{1}\mathbf{a}+z_{2}\mathbf{b}\|_{2}\geq\|\mathbf{u}\|_{2}$. Clearly, if $z_{1}z_{2}=0$ or $z_{1}=z_{2}$, we can easily deduce that $\|\mathbf{v}\|_{2}\geq\min\{\|\mathbf{a}\|_{2},\|\mathbf{b}\|_{2},\|\mathbf{a}+\mathbf{b}\|_{2}\}\geq\|\mathbf{u}\|_{2}$; Without loss of generality, we now assume $\|\mathbf{a}\|_{2}\leq\|\mathbf{b}\|_{2}$, $a_{1}b_{1}\geq 0,a_{2}b_{2}\leq 0$ and $z_{1}z_{2}\neq 0\wedge z_{1}\neq z_{2}$.
(1) $\mathbf{u}=\mathbf{a}$. In this case, we need to prove
$\|z_{1}\mathbf{a}+z_{2}\mathbf{b}\|_{2}\geq\|\mathbf{a}\|_{2}\Longleftrightarrow(z_{1}^{2}-1)(a_{1}^{2}+a_{2}^{2})+z_{2}^{2}(b_{1}^{2}+b_{2}^{2})+2z_{1}z_{2}(a_{1}b_{1}+a_{2}b_{2})\geq 0.$ (4)
If $(z_{1}z_{2})(a_{1}b_{1}+a_{2}b_{2})\geq 0$, the above inequality clearly holds. Meanwhile, from $\|\mathbf{a}\|_{2}\leq\|\mathbf{b}\|_{2}$ and $\|\mathbf{a}\|_{2}\leq\|\mathbf{a}+\mathbf{b}\|_{2},\|\mathbf{a}-\mathbf{b}\|_{2}$, we have
$a_{1}^{2}+a_{2}^{2}\leq b_{1}^{2}+b_{2}^{2}\ \wedge\ -(b_{1}^{2}+b_{2}^{2})\leq 2(a_{1}b_{1}+a_{2}b_{2})\leq b_{1}^{2}+b_{2}^{2}.$ (5)
Applying these bounds, we conclude that inequality (4) follows from
$(z_{1}^{2}-1)(a_{1}^{2}+a_{2}^{2})+z_{2}^{2}(b_{1}^{2}+b_{2}^{2})+2z_{1}z_{2}(a_{1}b_{1}+a_{2}b_{2})$
$\geq$ $(z_{1}^{2}-1)(a_{1}^{2}+a_{2}^{2})+z_{2}^{2}(b_{1}^{2}+b_{2}^{2})-z_{1}z_{2}(b_{1}^{2}+b_{2}^{2})$
$\geq$ $(z_{1}^{2}-1)(a_{1}^{2}+a_{2}^{2})+(z_{2}^{2}-z_{1}z_{2})(b_{1}^{2}+b_{2}^{2})\geq 0$
Thus, it remains to consider the case where $(z_{1}z_{2})(a_{1}b_{1}+a_{2}b_{2})<0$ and $|z_{1}|>|z_{2}|\geq 1$. We proceed by analyzing this scenario in detail.
  • $z_{1}z_{2}>0$ and $a_{1}b_{1}+a_{2}b_{2}<0$, we have
    $(z_{1}^{2}-1)(a_{1}^{2}+a_{2}^{2})+z_{2}^{2}(b_{1}^{2}+b_{2}^{2})+2z_{1}z_{2}(a_{1}b_{1}+a_{2}b_{2})$
    $=$ $(z_{1}^{2}-1)a_{1}^{2}+(z_{1}^{2}-1)a_{2}^{2}+z_{2}^{2}b_{1}^{2}+z_{2}^{2}b_{2}^{2}-2|z_{1}||z_{2}|(|a_{2}b_{2}+a_{1}b_{1}|)$
    $=$ $(z_{1}^{2}-1)a_{1}^{2}+(z_{1}^{2}-1)a_{2}^{2}+z_{2}^{2}b_{1}^{2}+z_{2}^{2}b_{2}^{2}-2|z_{1}||z_{2}|(|a_{2}||b_{2}|-a_{1}b_{1})$
    $=$ $z_{1}^{2}a_{2}^{2}+z_{2}^{2}b_{2}^{2}-2|z_{1}||z_{2}||a_{2}||b_{2}|+(z_{1}^{2}-1)a_{1}^{2}+z_{2}^{2}b_{1}^{2}+2z_{1}z_{2}a_{1}b_{1}-a_{2}^{2}$ (6)
    Clearly, if $|a_{1}|\geq|a_{2}|$, then the equation (6$)\geq(z_{1}^{2}-1)a_{1}^{2}-a_{2}^{2}\geq 3a_{1}^{2}-a_{2}^{2}\geq 0$, and, if $|b_{1}|\geq|a_{2}|$, then the equation (6$)\geq z_{2}^{2}b_{1}^{2}-a_{2}^{2}\geq b_{1}^{2}-a_{2}^{2}\geq 0$. Consequently, the remaining case to consider is $|a_{1}|<|a_{2}|\wedge|b_{1}|<|a_{2}|$. From $|a_{1}|<|a_{2}|$ and the definition of the reduced base, we know $|b_{1}|\geq|b_{2}|$. Combining this with $|b_{1}|<|a_{2}|$ and the bound $a_{1}^{2}+a_{2}^{2}\leq b_{1}^{2}+b_{2}^{2}$ (equation (5)), we deduce $|a_{1}|<|b_{2}|$. Thus, the remaining case is
    $|a_{1}|<|b_{2}|\leq|b_{1}|<|a_{2}|.$
    In this case, if $|z_{2}|\geq 2$, the the equation (6$)\geq z_{2}^{2}b_{1}^{2}-a_{2}^{2}\geq 4b_{1}^{2}-a_{2}^{2}\geq 2b_{1}^{2}+2b_{2}^{2}-a_{2}^{2}\geq 2(a_{1}^{2}+a_{2}^{2})-a_{2}^{2}\geq 0.$ If $|z_{2}|=1$, the equation (6) is
    $(b_{1}^{2}+b_{2}^{2})-(a_{1}^{2}+a_{2}^{2})+2|z_{1}|a_{1}b_{1}+z_{1}^{2}(a_{1}^{2}+a_{2}^{2})-2|z_{1}||a_{2}||b_{2}|$
    $\geq$ $z_{1}^{2}a_{2}^{2}-2|z_{1}||a_{2}||b_{2}|\geq z_{1}^{2}|a_{2}||b_{2}|-2|z_{1}||a_{2}||b_{2}|=(z_{1}^{2}-2|z_{1}|)|a_{2}||b_{2}|\geq 0.$
  • $z_{1}z_{2}<0$ and $a_{1}b_{1}+a_{2}b_{2}>0$, we have
    $(z_{1}^{2}-1)(a_{1}^{2}+a_{2}^{2})+z_{2}^{2}(b_{1}^{2}+b_{2}^{2})+2z_{1}z_{2}(a_{1}b_{1}+a_{2}b_{2})$
    $=$ $(z_{1}^{2}-1)a_{1}^{2}+(z_{1}^{2}-1)a_{2}^{2}+z_{2}^{2}b_{1}^{2}+z_{2}^{2}b_{2}^{2}-2|z_{1}||z_{2}|(a_{1}b_{1}-|a_{2}b_{2}|)$
    $=$ $z_{1}^{2}a_{1}^{2}+z_{2}^{2}b_{1}^{2}-2|z_{1}||z_{2}|a_{1}b_{1}+(z_{1}^{2}-1)a_{2}^{2}+z_{2}^{2}b_{2}^{2}+2|z_{1}||z_{2}||a_{2}||b_{2}|-a_{1}^{2}.$ (7)
    Clearly, if $|a_{2}|\geq|a_{1}|$, then the equation (7$)\geq(z_{1}^{2}-1)a_{2}^{2}-a_{1}^{2}\geq 3a_{2}^{2}-a_{1}^{2}\geq 0$, and, if $|b_{2}|\geq|a_{1}|$, then the equation (7$)\geq z_{2}^{2}b_{2}^{2}-a_{1}^{2}\geq b_{2}^{2}-a_{1}^{2}\geq 0$. Consequently, the left case is $|a_{1}|>|a_{2}|\wedge|a_{1}|>|b_{2}|$. From $|a_{1}|>|a_{2}|$ and the definition of the reduced base, we know $|b_{1}|\leq|b_{2}|$, and, $|a_{1}|>|b_{2}|$ and $a_{1}^{2}+a_{2}^{2}\leq b_{1}^{2}+b_{2}^{2}$ (equation (5)) implies $|a_{2}|<|b_{1}|$. That is, the left case is
    $|a_{2}|<|b_{1}|\leq|b_{2}|<|a_{1}|.$
    In this case, if $|z_{2}|\geq 2$, then the equation (7$)\geq z_{2}^{2}b_{2}^{2}-a_{1}^{2}\geq 4b_{2}^{2}-a_{2}^{2}\geq 2b_{1}^{2}+2b_{2}^{2}-a_{1}^{2}\geq 2(a_{1}^{2}+a_{2}^{2})-a_{1}^{2}\geq 0.$ If $|z_{2}|=1$, the equation (7) is
    $z_{1}^{2}(a_{1}^{2}+a_{2}^{2})+(b_{1}^{2}+b_{2}^{2})-(a_{1}^{2}+a_{2}^{2})-2|z_{1}|a_{1}b_{1}+2|z_{1}||a_{2}||b_{2}|$
    $\geq$ $z_{1}^{2}a_{1}^{2}-2|z_{1}|a_{1}b_{1}\geq z_{1}^{2}|a_{1}||b_{1}|-2|z_{1}|a_{1}b_{1}=(z_{1}^{2}-2|z_{1}|)a_{1}b_{1}\geq 0.$
(2) $\mathbf{u}=\mathbf{a}+\mathbf{b}$. In this case, we need to prove
$||z_{1}\mathbf{a}+z_{2}\mathbf{b}||\geq||\mathbf{a}+\mathbf{b}||\Longleftrightarrow(z_{1}^{2}-1)(a_{1}^{2}+a_{2}^{2})+(z_{2}^{2}-1)(b_{1}^{2}+b_{2}^{2})+2(z_{1}z_{2}-1)(a_{1}b_{1}+a_{2}b_{2})\geq 0.$ (8)
From $\|\mathbf{a}+\mathbf{b}\|_{2}\leq\|\mathbf{a}\|_{2}\leq\|\mathbf{b}\|_{2}$, we have
$a_{1}b_{1}+a_{2}b_{2}<0\ \wedge\ a_{1}^{2}+a_{2}^{2}\leq b_{1}^{2}+b_{2}^{2}\leq-2(a_{1}b_{1}+a_{2}b_{2})=2(|a_{2}||b_{2}|-|a_{1}||b_{1}|)\leq(a_{1}^{2}+a_{2}^{2})+(b_{1}^{2}+b_{2}^{2}).$ (9)
Since $a_{1}b_{1}+a_{2}b_{2}<0$, the target inequality (8) clearly holds if $z_{1}z_{2}-1\leq 0$. We only need to consider the case that $z_{1}z_{2}>1$.
In fact, by equation (9),
$(z_{1}^{2}-1)(a_{1}^{2}+a_{2}^{2})+(z_{2}^{2}-1)(b_{1}^{2}+b_{2}^{2})+2(z_{1}z_{2}-1)(a_{1}b_{1}+a_{2}b_{2})$
$=$ $(z_{1}^{2}-1)(a_{1}^{2}+a_{2}^{2})+(z_{2}^{2}-1)(b_{1}^{2}+b_{2}^{2})+2z_{1}z_{2}(a_{1}b_{1}+a_{2}b_{2})-2(a_{1}b_{1}+a_{2}b_{2})$
$=$ $z_{1}^{2}a_{1}^{2}+z_{1}^{2}a_{2}^{2}-(a_{1}^{2}+a_{2}^{2})+z_{2}^{2}b_{1}^{2}+z_{2}^{2}b_{2}^{2}-(b_{1}^{2}+b_{2}^{2})-2z_{1}z_{2}|a_{2}||b_{2}|+2z_{1}z_{2}|a_{1}||b_{1}|-2(a_{1}b_{1}+a_{2}b_{2})$
$=$ $(z_{1}^{2}a_{2}^{2}-2z_{1}z_{2}|a_{2}||b_{2}|+z_{2}^{2}b_{2}^{2})+2z_{1}z_{2}|a_{1}||b_{1}|+(z_{1}^{2}-1)a_{1}^{2}+z_{2}^{2}b_{1}^{2}+(-2(a_{1}b_{1}+a_{2}b_{2})-(b_{1}^{2}+b_{2}^{2}))-a_{2}^{2}$ (10)
$=$ $(z_{1}^{2}a_{2}^{2}-2z_{1}z_{2}|a_{2}||b_{2}|+z_{2}^{2}b_{2}^{2})+2z_{1}z_{2}|a_{1}||b_{1}|+z_{1}^{2}a_{1}^{2}+(z_{2}^{2}-1)b_{1}^{2}+(-2(a_{1}b_{1}+a_{2}b_{2})-(a_{1}^{2}+a_{2}^{2}))-b_{2}^{2}$ (11)
If $|b_{1}|\geq|a_{2}|$, then the equation (10$)\geq z_{2}^{2}b_{1}^{2}-a_{2}^{2}\geq b_{1}^{2}-a_{2}^{2}\geq 0$. Hence, we only argue the case that $|b_{1}|<|a_{2}|$. Here, since $a_{1}^{2}+a_{2}^{2}\leq b_{1}^{2}+b_{2}^{2}$, we have $|b_{2}|>|a_{1}|$. Also, $(|a_{1}|-|a_{2}|)(|b_{1}|-|b_{2}|)\leq 0$ implies that (1) $|b_{1}|<|a_{2}|\leq|a_{1}|<|b_{2}|$ or (2) $|a_{1}|<|b_{2}|\leq|b_{1}|<|a_{2}|$. Now, we argue each case as follows.
  • $|b_{1}|<|a_{2}|\leq|a_{1}|<|b_{2}|$. If $|z_{1}|\neq 1$, then the equation (10$)\geq(z_{1}^{2}-1)a_{1}^{2}-a_{2}^{2}\geq 3a_{1}^{2}-a_{2}^{2}\geq 0$. If $|z_{1}|=1$, then $|z_{2}|\geq 2$ and the equation (10) is
    $z_{2}^{2}b_{1}^{2}+z_{2}^{2}b_{2}^{2}-2|z_{2}||a_{2}||b_{2}|+2|z_{2}||a_{1}||b_{1}|+(2(|a_{2}||b_{2}|-|a_{1}||b_{1}|)-(b_{1}^{2}+b_{2}^{2}))$
    $\geq$ $z_{2}^{2}b_{2}^{2}-2|z_{2}||a_{2}||b_{2}|\geq z_{2}^{2}|a_{2}||b_{2}|-2|z_{2}||a_{2}||b_{2}|=(z_{2}^{2}-2|z_{2}|)|a_{2}||b_{2}|\geq 0.$
[…]
    $z_{1}^{2}a_{2}^{2}-2|z_{1}||a_{2}||b_{2}|+2|z_{1}||a_{1}||b_{1}|+z_{1}^{2}a_{1}^{2}+(-2(a_{1}b_{1}+a_{2}b_{2})-(a_{1}^{2}+a_{2}^{2}))$
    $\geq$ $z_{1}^{2}a_{2}^{2}-2|z_{1}||a_{2}||b_{2}|\geq z_{1}^{2}|a_{2}||b_{2}|-2|z_{1}||a_{2}||b_{2}|\geq(z_{1}^{2}-2|z_{1}|)|a_{2}||b_{2}|\geq 0.$
(3) $\mathbf{u}=\mathbf{a}-\mathbf{b}$. The proof for this case follows the same logic as case $\mathbf{u}=\mathbf{a}+\mathbf{b}$, and thus we omit the details.
Combining the three cases (1)-(3) above, we complete the proof of $\lambda_{1}(\mathcal{L})$. To prove $\lambda_{2}(\mathcal{L})=\|\mathbf{v}\|_{2}$, by Lemma 2, it suffices to show that the pair $[\mathbf{u},\mathbf{v}]$ is Lagrange-reduced, i.e.,
$\|\mathbf{u}\|_{2}\leq\|\mathbf{v}\|_{2}\leq\min\left\{\|\mathbf{u}+\mathbf{v}\|_{2},\|\mathbf{u}-\mathbf{v}\|_{2}\right\}.$
In fact, by the choice of $q$, we have:
$\|\mathbf{v}\|_{2}$ $=\|\mathbf{x}-q\mathbf{u}\|_{2}\leq\|\mathbf{x}-(q+1)\mathbf{u}\|_{2}=\|\mathbf{v}-\mathbf{u}\|_{2},$
$\|\mathbf{v}\|_{2}$ $=\|\mathbf{x}-q\mathbf{u}\|_{2}\leq\|\mathbf{x}-(q-1)\mathbf{u}\|_{2}=\|\mathbf{v}+\mathbf{u}\|_{2}.$
Thus, the inequalities $\|\mathbf{v}\|_{2}\leq\|\mathbf{u}\pm\mathbf{v}\|_{2}$ hold, completing the proof.
IV Our new reduced algorithm and its analysis
Building upon our new defined reduced basis (Definition 7) and its properties (Theorem 1), we introduce an algorithm that directly transforms any given two-dimensional lattice basis into a reduced basis. This approach eliminates the necessity for an initial conversion to HNF as employed by Eisenbrand [4].
IV-A Design Idea
For any input lattice basis $[\mathbf{a}\ \mathbf{b}]$ with $\mathbf{a}=(a_{1}\ a_{2})^{T}$, $\mathbf{b}=(b_{1}\ b_{2})^{T}$ and $\|\mathbf{a}\|\geq\|\mathbf{b}\|$, if it is not reduced, then we have two possible cases: (1) $a_{1}a_{2}b_{1}b_{2}>0$ or (2) $a_{1}a_{2}b_{1}b_{2}\leq 0$ and $(|a_{1}|-|a_{2}|)(|b_{1}|-|b_{2}|)>0$.
We first consider the case (2), without loss of generality, assume $a_{1}\geq b_{1}\geq 0$ and $a_{2}b_{2}\leq 0$. If these conditions are not met, we can adjust the basis vectors using negation or swapping operations to achieve them.
If $|a_{1}|>|a_{2}|$ and $|b_{1}|>|b_{2}|$, inspired by prior work [4, 3], we can compute a new basis $[\mathbf{b}\ \mathbf{c}]$ with $\mathbf{c}=\mathbf{a}-q\mathbf{b}$, where $q=\lfloor\frac{a_{1}}{b_{1}}\rfloor_{o}>0$. Then
$a_{1}>b_{1}>c_{1}\geq 0,\ b_{2}c_{2}\leq 0,\ |c_{2}|=|a_{2}-qb_{2}|=|a_{2}|+q|b_{2}|>|b_{2}|.$
As a result, $b_{1}b_{2}c_{1}c_{2}\leq 0$. If $|c_{1}|-|c_{2}|\leq 0$, the new basis $[\mathbf{b}\ \mathbf{c}]$ is reduced. Otherwise, we have $|b_{1}|-|b_{2}|>|c_{1}|-|c_{2}|>0$. Repeating this process iteratively reduces the difference between the absolute values of the first and second coordinates of the new vector until the terminal condition (the equation (2)) is satisfied.
If $|a_{1}|<|a_{2}|$ and $|b_{1}|<|b_{2}|$, in this case, we focus on the second coordinate. Using a similar process, we compute a new basis $[\mathbf{b}\ \mathbf{c}]$, where $\mathbf{c}=\mathbf{a}-q\mathbf{b}$ and $q=\lfloor\frac{a_{2}}{b_{2}}\rfloor_{o}\leq 0$. Then
$c_{1}>a_{1}>b_{1}\geq 0,\ b_{2}c_{2}\leq 0,\ |c_{2}|=|a_{2}-qb_{2}|=|a_{2}|-|q||b_{2}|<|b_{2}|.$
Thus, $b_{1}b_{2}c_{1}c_{2}\leq 0$. If $|c_{1}|-|c_{2}|\geq 0$, the new basis $[\mathbf{b}\ \mathbf{c}]$ is reduced. Otherwise, we have $|b_{1}|-|b_{2}|<|c_{1}|-|c_{2}|<0$. Repeating this process iteratively reduces the difference between the absolute values of the coordinates until the terminal condition (the equation (2)) is satisfied.
For the case (1). When $a_{1}a_{2}b_{1}b_{2}>0$, the key is to transform the basis such that $a_{1}a_{2}b_{1}b_{2}\leq 0$. Since $\mathbf{a}$ and $\mathbf{b}$ are linearly independent, it holds that $\frac{a_{1}}{b_{1}}\neq\frac{a_{2}}{b_{2}}$. Let $q_{1}\triangleq\lfloor\frac{a_{1}}{b_{1}}\rfloor_{o},q_{2}\triangleq\lfloor\frac{a_{2}}{b_{2}}\rfloor_{o}$. If $q_{1}\neq q_{2}$, we choose
$q=\begin{cases}q_{1},&\text{if }|a_{1}|\geq|a_{2}|\wedge q_{1}\geq q_{2},\\ q_{1}+1,&\text{if }|a_{1}|\geq|a_{2}|\wedge q_{1}<q_{2},\\ q_{2},&\text{if }|a_{1}|<|a_{2}|\wedge q_{2}\geq q_{1},\\ q_{2}+1,&\text{if }|a_{1}|<|a_{2}|\wedge q_{2}<q_{1}.\end{cases}$
Then the new basis $[\mathbf{b}\ \mathbf{c}]$ with $\mathbf{c}=\mathbf{a}-q\mathbf{b}$ ensures $b_{1}b_{2}c_{1}c_{2}\leq 0$, achieving the goal. Otherwise $q=q_{1}=q_{2}$, iteratively reduce the basis $[\mathbf{b}\ \mathbf{c}]$ with the same method until $q_{1}\neq q_{2}$.
IV-B Algorithm Description and Analysis
Based on the proposed design concept, the detailed steps of our algorithm are provided in Algorithm 3. It should be noted that the Algorithm 3 invokes two sub-algorithms: UMTrans1 and UMTrans2, which are variants of UMTrans1 (Algorithm 1) and UMTrans2 (Algorithm 2), respectively. The only difference is whether the returned result includes the integer $q$.
Before discussing the correctness and complexity of Algorithm 3, we first present two lemmas to analyze the properties of Algorithm 1 and Algorithm 2.
Algorithm 1 UMTrans1($\mathbf{a},\mathbf{b}$)
0:  A basis $[\mathbf{a}\ \mathbf{b}]$ where $\mathbf{a}=(a_{1}\ a_{2})^{T}$ and $\mathbf{b}=(b_{1}\ b_{2})^{T}$, satisfying $a_{1}a_{2}b_{1}b_{2}>0$
0:  A new base $\mathbf{a}=(a_{1}\ a_{2})^{T}$, $\mathbf{b}=(b_{1}\ b_{2})^{T}$ and an integer $q$
1:  if ($|a_{1}|\geq|a_{2}|$)
2:   $q:=\left\lfloor\frac{a_{1}}{b_{1}}\right\rfloor_{o}$
3:   if ($|a_{2}-qb_{2}|\geq|b_{2}|\wedge sgn(a_{2})=sgn(a_{2}-qb_{2})$), $q:=q+1$
4:  else
5:   $q:=\left\lfloor\frac{a_{2}}{b_{2}}\right\rfloor_{o}$
6:   if ($|a_{1}-qb_{1}|\geq|b_{1}|\wedge sgn(a_{1})=sgn(a_{1}-qb_{1})$), $q:=q+1$
7:  ($\mathbf{a},\mathbf{b}$) := ($\mathbf{b},\mathbf{a}-q\mathbf{b}$)
8:  return $\mathbf{a},\mathbf{b},q$
Algorithm 2 UMTrans2($\mathbf{a},\mathbf{b}$)
0:  A basis $[\mathbf{a}\ \mathbf{b}]$ where $\mathbf{a}=(a_{1}\ a_{2})^{T}$ and $\mathbf{b}=(b_{1}\ b_{2})^{T}$, satisfying $a_{1}a_{2}b_{1}b_{2}\leq 0$ and $(|a_{1}|-|a_{2}|)(|b_{1}|-|b_{2}|)>0$
0:  A new basis $\mathbf{a}=(a_{1}\ a_{2})^{T}$, $\mathbf{b}=(b_{1}\ b_{2})^{T}$ and an integer $q$
1:  if ($|a_{1}|>|a_{2}|$)
2:   $q:=\left\lfloor\frac{a_{1}}{b_{1}}\right\rfloor_{o}$
3:  else
4:   $q:=\left\lfloor\frac{a_{2}}{b_{2}}\right\rfloor_{o}$
5:  ($\mathbf{a},\mathbf{b}$) := ($\mathbf{b},\mathbf{a}-q\mathbf{b}$)
6:  return $\mathbf{a},\mathbf{b},q$
Algorithm 3 CrossEuc($\mathbf{a},\mathbf{b}$)
0:  A basis $[\mathbf{a}\ \mathbf{b}]\in\mathbb{Z}^{2\times 2}$ with $\mathbf{a}=(a_{1}\ a_{2})^{T}$, $\mathbf{b}=(b_{1}\ b_{2})^{T}$
0:  A new basis $\mathbf{a}=(a_{1}\ a_{2})^{T}$, $\mathbf{b}=(b_{1}\ b_{2})^{T}$ satisfy $a_{1}a_{2}b_{1}b_{2}\leq 0\wedge(|a_{1}|-|a_{2}|)(|b_{1}|-|b_{2}|)\leq 0$
1:  While ($a_{1}a_{2}b_{1}b_{2}>0$), do
2:   $(\mathbf{a},\mathbf{b})\leftarrow\textbf{UMTrans1}^{\star}(\mathbf{a},\mathbf{b})$
3:  Endwhile
4:  While ($(|a_{1}|-|a_{2}|)(|b_{1}|-|b_{2}|)>0$), do
5:   $(\mathbf{a},\mathbf{b})\leftarrow\textbf{UMTrans2}^{\star}(\mathbf{a},\mathbf{b})$
6:  Endwhile
7:  If (($a_{1}b_{1}=0\wedge sgn(a_{2})\neq sgn(b_{2}))$ or ($a_{2}b_{2}=0\wedge sgn(a_{1})\neq sng(b_{1})$)), $\mathbf{b}:=-\mathbf{b}$.
8:  If $\|\mathbf{a}\|>\|\mathbf{b}\|$, Swap($\mathbf{a},\mathbf{b}$)
9:  If $\mathbf{a}=[0,0]$, Return[$\mathbf{a},\mathbf{b}$]
10:  Else
11:   if ($a_{2}b_{2}\leq 0$)
12:    $\mathbf{b}:=\min\left\{\mathbf{b}-\left\lfloor\frac{|b_{1}|-|b_{2}|}{|a_{1}|+|a_{2}|}\right\rfloor\mathbf{a},\mathbf{b}-\left\lceil\frac{|b_{1}|-|b_{2}|}{|a_{1}|+|a_{2}|}\right\rceil\mathbf{a}\right\}$
13:   else
14:    $\mathbf{b}:=\min\left\{\mathbf{b}-\left\lfloor\frac{|b_{2}|-|b_{1}|}{|a_{1}|+|a_{2}|}\right\rfloor\mathbf{a},\mathbf{b}-\left\lceil\frac{|b_{2}|-|b_{1}|}{|a_{1}|+|a_{2}|}\right\rceil\mathbf{a}\right\}$
15:   Return $[\mathbf{a}\ \mathbf{b}]$.
Lemma 5.
For any lattice basis $\mathbf{B}$ defined in equation (1) satisfying $\|\mathbf{a}\|\geq\|\mathbf{b}\|$ and $a_{1}a_{2}b_{1}b_{2}>0$, let $q_{i}=\lfloor\frac{a_{i}}{b_{i}}\rfloor_{o}$ for $i=1,2$.
If $|a_{1}|\geq|a_{2}|$, define
$q:=\begin{cases}q_{1},&\text{if }q_{1}\geq q_{2},\\ q_{1}+1,&\text{if }q_{1}<q_{2}.\end{cases}$
If $|a_{1}|<|a_{2}|$, define
$q:=\begin{cases}q_{2},&\text{if }q_{2}\geq q_{1},\\ q_{2}+1,&\text{if }q_{2}<q_{1}.\end{cases}$
Then the new matrix
$\mathbf{B}^{\prime}=[\mathbf{a}^{\prime}\ \mathbf{b}^{\prime}]=\begin{pmatrix}a_{1}^{\prime}&b_{1}^{\prime}\\ a_{2}^{\prime}&b_{2}^{\prime}\end{pmatrix}:=[\mathbf{b}\ \mathbf{a}-q\mathbf{b}]=\mathbf{B}\begin{pmatrix}0&1\\ 1&-q\end{pmatrix}\in\mathbb{R}^{2\times 2}$
forms a valid basis, satisfying one of the following conditions:
  1. 1.
    $a_{1}^{\prime}a_{2}^{\prime}b_{1}^{\prime}b_{2}^{\prime}\leq 0$, with $(\|\mathbf{a}^{\prime}\|=\|\mathbf{b}\|)\wedge(\|\mathbf{b}^{\prime}\|<\max\{\|\mathbf{a}\|\cdot\|\mathbf{b}\|,\|\mathbf{a}\|,\|\mathbf{b}\|\})$ or
  2. 2.
    $a_{1}^{\prime}a_{2}^{\prime}b_{1}^{\prime}b_{2}^{\prime}>0$, with $(\|\mathbf{b}^{\prime}\|\leq\min\{\|\mathbf{a}-\mathbf{b}\|,\|\mathbf{a}+\mathbf{b}\|\})\wedge(\|\mathbf{b}^{\prime}\|<\|\mathbf{a}^{\prime}\|)\wedge(\|\mathbf{b}^{\prime}\|<\frac{1}{2}\|\mathbf{a}\|)$.
Proof.
Due to the property of the unimodular matrix transformation, it is evident that $\mathbf{B}^{\prime}$ forms a basis. Without loss of generality, we assume $a_{1}b_{1}>0$ and $a_{2}b_{2}>0$, focusing on the case where $|a_{1}|\geq|a_{2}|$; the argument for $|a_{1}|<|a_{2}|$ follows analogously.
Let $a_{1}=q_{1}b_{1}+r_{1}$ and $a_{2}=q_{2}b_{2}+r_{2}$. Then $\text{sgn}(r_{1})=\text{sgn}(a_{1})=\text{sgn}(b_{1})$, $\text{sgn}(r_{2})=\text{sgn}(a_{2})=\text{sgn}(b_{2})$, and $0\leq|r_{1}|<|b_{1}|,0\leq|r_{2}|<|b_{2}|.$
(1) $q_{1}>q_{2}.$ In this case,
$\begin{pmatrix}a_{1}^{\prime}&b_{1}^{\prime}\\ a_{2}^{\prime}&b_{2}^{\prime}\end{pmatrix}:=\begin{pmatrix}b_{1}&a_{1}-q_{1}b_{1}\\ b_{2}&a_{2}-q_{1}b_{2}\end{pmatrix}.$
Clearly, $\text{sgn}(a_{1}^{\prime})=\text{sgn}(b_{1})=\text{sgn}(a_{1}-q_{1}b_{1})=\text{sgn}(b_{1}^{\prime})$. Assuming $q_{1}=q_{2}+k$ with $1\leq k\leq q_{1}=\lfloor\frac{a_{1}}{b_{1}}\rfloor_{o}\leq\frac{a_{1}}{b_{1}}$, we have
$0\leq|a_{2}-q_{2}b_{2}|=|r_{2}|<|b_{2}|$
$\implies$ $(k-1)|b_{2}|<|a_{2}-q_{1}b_{2}|=|r_{2}-kb_{2}|=k|b_{2}|-|r_{2}|\leq k|b_{2}|\leq\frac{a_{1}}{b_{1}}|b_{2}|\leq|a_{1}||b_{2}|\leq\|\mathbf{a}\|\cdot\|\mathbf{b}\|$
$\implies$ $\|\mathbf{b}^{\prime}\|=\max\{|a_{1}-q_{1}b_{1}|,|a_{2}-q_{1}b_{2}|\}=\max\{|r_{1}|,|r_{2}-kb_{2}|\}\leq\|\mathbf{a}\|\cdot\|\mathbf{b}\|$
and $\text{sgn}(a_{2}^{\prime})=\text{sgn}(b_{2})\neq\text{sgn}(a_{2}-q_{1}b_{2})=\text{sgn}(b_{2}^{\prime})$, which implies $a_{1}^{\prime}a_{2}^{\prime}b_{1}^{\prime}b_{2}^{\prime}\leq 0.$
(2) $q_{1}<q_{2}.$ In this case,
$\begin{pmatrix}a_{1}^{\prime}&b_{1}^{\prime}\\ a_{2}^{\prime}&b_{2}^{\prime}\end{pmatrix}:=\begin{pmatrix}b_{1}&a_{1}-(q_{1}+1)b_{1}\\ b_{2}&a_{2}-(q_{1}+1)b_{2}\end{pmatrix}.$
Clearly, $|b_{1}^{\prime}|=|a_{1}-(q_{1}+1)b_{1}|=|r_{1}-b_{1}|\leq|b_{1}|=|a_{1}^{\prime}|,$ and $\text{sgn}(a_{1}^{\prime})=\text{sgn}(b_{1})\neq\text{sgn}(a_{1}-(q_{1}+1)b_{1})=\text{sgn}(b_{1}^{\prime})$. Assuming $q_{2}=q_{1}+k$ with $k\geq 1$, we have
$|a_{2}-(q_{1}+1)b_{2}|=|a_{2}-(q_{2}-k+1)b_{2}|=|r_{2}+(k-1)b_{2}|<k|b_{2}|<q_{2}|b_{2}|<|a_{2}|\leq\|\mathbf{a}\|$
$\implies$ $\|\mathbf{b}^{\prime}\|=\max\{|b_{1}^{\prime}|,|b_{2}^{\prime}|\}=\max\{|r_{1}-b_{1}|,|r_{2}+(k-1)b_{2}|\}\leq\max\{|b_{1}|,|a_{2}|\}\leq\max\{\|\mathbf{a}\|,\|\mathbf{b}\|\},$
and $\text{sgn}(a_{2}^{\prime})=\text{sgn}(b_{2})=\text{sgn}(a_{2}-(q_{1}+1)b_{2})=\text{sgn}(b_{2}^{\prime})$, leading to $a_{1}^{\prime}a_{2}^{\prime}b_{1}^{\prime}b_{2}^{\prime}\leq 0.$
(3) $q_{1}=q_{2}.$ In this case,
$\begin{pmatrix}a_{1}^{\prime}&b_{1}^{\prime}\\ a_{2}^{\prime}&b_{2}^{\prime}\end{pmatrix}:=\begin{pmatrix}b_{1}&a_{1}-q_{1}b_{1}\\ b_{2}&a_{2}-q_{1}b_{2}\end{pmatrix}=\begin{pmatrix}b_{1}&r_{1}\\ b_{2}&r_{2}\end{pmatrix}.$
If $r_{1}=0$ or $r_{2}=0$, then $a_{1}^{\prime}a_{2}^{\prime}b_{1}^{\prime}b_{2}^{\prime}\leq 0$ and $\max\{|a_{1}^{\prime}|,|a_{2}^{\prime}|,|b_{1}^{\prime}|,|b_{2}^{\prime}|\}=\max\{|b_{1}|,|b_{2}|\}$. Otherwise, $a_{1}^{\prime}a_{2}^{\prime}b_{1}^{\prime}b_{2}^{\prime}>0$ and
$\left.\begin{array}[]{c}(|b_{1}^{\prime}|=|r_{1}|=|a_{1}-q_{1}b_{1}|\leq\min\{|a_{1}-b_{1}|,|a_{1}+b_{1}|\})\wedge(|b_{1}^{\prime}|=|r_{1}|<|b_{1}|=|a_{1}^{\prime}|)\wedge(|b_{1}^{\prime}|=|r_{1}|<\frac{|a_{1}|}{2}),\\ (|b_{2}^{\prime}|=|r_{2}|=|a_{2}-q_{2}b_{2}|\leq\min\{|a_{2}-b_{2}|,|a_{2}+b_{2}|\})\wedge(|b_{2}^{\prime}|=|r_{2}|<|b_{2}|=|a_{2}^{\prime}|)\wedge(|b_{2}^{\prime}|=|r_{2}|<\frac{|a_{2}|}{2}).\end{array}\right\}$
$\implies(\|\mathbf{b}^{\prime}\|\leq\min\{\|\mathbf{a}-\mathbf{b}\|,\|\mathbf{a}+\mathbf{b}\|\})\wedge(\|\mathbf{b}^{\prime}\|<\|\mathbf{a}^{\prime}\|)\wedge(\|\mathbf{b}^{\prime}\|<\frac{1}{2}\|\mathbf{a}\|).$
Lemma 6.
Let $\mathbf{B}$ be a lattice basis as defined in equation (1) satisfying $\|\mathbf{a}\|\geq\|\mathbf{b}\|$, $a_{1}a_{2}b_{1}b_{2}\leq 0$ and $(\left|a_{1}\right|-\left|a_{2}\right|)(\left|b_{1}\right|-\left|b_{2}\right|)>0$.
If $|a_{1}|>|a_{2}|$, define $q_{1}=\lfloor\frac{a_{1}}{b_{1}}\rfloor_{o}$ and
$\mathbf{B}^{\prime}=[\mathbf{a}^{\prime}\ \mathbf{b}^{\prime}]=\begin{pmatrix}a_{1}^{\prime}&b_{1}^{\prime}\\ a_{2}^{\prime}&b_{2}^{\prime}\end{pmatrix}:=\begin{pmatrix}b_{1}&a_{1}-q_{1}b_{1}\\ b_{2}&a_{2}-q_{1}b_{2}\end{pmatrix}=[\mathbf{b}\ \mathbf{a}-q_{1}\mathbf{b}]=\mathbf{B}\begin{pmatrix}0&1\\ 1&-q_{1}\end{pmatrix}$
If $|a_{1}|<|a_{2}|$, define $q_{2}=\lfloor\frac{a_{2}}{b_{2}}\rfloor_{o}$ and
$\mathbf{B}^{\prime}=[\mathbf{a}^{\prime}\ \mathbf{b}^{\prime}]=\begin{pmatrix}a_{1}^{\prime}&b_{1}^{\prime}\\ a_{2}^{\prime}&b_{2}^{\prime}\end{pmatrix}:=\begin{pmatrix}b_{1}&a_{1}-q_{2}b_{1}\\ b_{2}&a_{2}-q_{2}b_{2}\end{pmatrix}=[\mathbf{b}\ \mathbf{a}-q_{2}\mathbf{b}]=\mathbf{B}\begin{pmatrix}0&1\\ 1&-q_{2}\end{pmatrix}$
Then the new matrix $\mathbf{B}^{\prime}$ forms a valid basis, satisfying one of the following conditions:
  1. 1.
    $\mathbf{B}^{\prime}$ is reduced, or
  2. 2.
    $a_{1}^{\prime}a_{2}^{\prime}b_{1}^{\prime}b_{2}^{\prime}\leq 0$, with $(\|\mathbf{b}^{\prime}\|\leq\min\{\|\mathbf{a}-\mathbf{b}\|,\|\mathbf{a}+\mathbf{b}\|\})\wedge(\|\mathbf{b}^{\prime}\|<\|\mathbf{a}^{\prime}\|)\wedge(\|\mathbf{b}^{\prime}\|<\frac{1}{2}\|\mathbf{a}\|)\wedge(||b_{1}^{\prime}|-|b_{2}^{\prime}||<\frac{||a_{1}|-|a_{2}||}{2})$.
Proof.
Clearly, $\mathbf{B}^{\prime}$ forms a basis as a direct consequence of the properties of unimodular matrix transformations. Without loss of generality, we focus on the case where $|a_{1}|>|a_{2}|$; the argument for $|a_{1}|<|a_{2}|$ follows analogously.
In case that $|a_{1}|>|a_{2}|$, since $a_{1}a_{2}b_{1}b_{2}\leq 0$ and $(|a_{1}|-|a_{2}|)(|b_{1}|-|b_{2}|)>0$, we have $|a_{1}|>|a_{2}|\geq 0$, $|b_{1}|>|b_{2}|\geq 0$ and, without loss of generality, we assume $a_{1}b_{1}>0$ and $a_{2}b_{2}\leq 0$. Let $a_{1}=q_{1}b_{1}+r_{1}$. Then $q_{1}>0$, $\text{sgn}(r_{1})=\text{sgn}(a_{1})=\text{sgn}(b_{1})$, and
$0\leq|r_{1}|<|b_{1}|\implies 0\leq|r_{1}|<\frac{|a_{1}|}{2}.$
Additionally, $\text{sgn}(a_{1}^{\prime})=\text{sgn}(b_{1})=\text{sgn}(a_{1}-q_{1}b_{1})=\text{sgn}(b_{1}^{\prime})$ and $\text{sgn}(a_{2}^{\prime})=\text{sgn}(b_{2})\neq\text{sgn}(a_{2}-q_{1}b_{2})=\text{sgn}(b_{2}^{\prime})$, which implies $a_{1}^{\prime}a_{2}^{\prime}b_{1}^{\prime}b_{2}^{\prime}\leq 0.$ Furthermore, if $|b_{1}^{\prime}|-|b_{2}^{\prime}|\leq 0$, then $(|a_{1}^{\prime}|-|a_{2}^{\prime}|)(|b_{1}^{\prime}|-|b_{2}^{\prime}|)\leq 0$ and thus $\mathbf{B}^{\prime}$ is reduced. Otherwise, $|b_{2}^{\prime}|<|b_{1}^{\prime}|=|r_{1}|$, which implies
$\|\mathbf{b}^{\prime}\|=\max\{|b_{1}^{\prime}|,|b_{2}^{\prime}|\}=|b_{1}^{\prime}|=|r_{1}|=|a_{1}-q_{1}b_{1}|\leq\min\{|a_{1}-b_{1}|,|a_{1}+b_{1}|\}\leq\min\{\|\mathbf{a}-\mathbf{b}\|,\|\mathbf{a}+\mathbf{b}\|\},$
$\|\mathbf{b}^{\prime}\|=\max\{|b_{1}^{\prime}|,|b_{2}^{\prime}|\}=|b_{1}^{\prime}|=|r_{1}|<|b_{1}|\leq\max\{|b_{1}|,|b_{2}|\}=\|\mathbf{b}\|=\|\mathbf{a}^{\prime}\|,$
$\|\mathbf{b}^{\prime}\|=\max\{|b_{1}^{\prime}|,|b_{2}^{\prime}|\}=|b_{1}^{\prime}|=|r_{1}|<\frac{|a_{1}|}{2}=\frac{1}{2}\max\{|a_{1}|,|a_{2}|\}=\frac{1}{2}\|\mathbf{a}\|,$
$|b_{1}^{\prime}|-|b_{2}^{\prime}|=|r_{1}|-|a_{2}-q_{1}b_{2}|=|r_{1}|-(|a_{2}|+q_{1}|b_{2}|)\leq|r_{1}|-|a_{2}|<\frac{|a_{1}|-|a_{2}|}{2}.$
Based on Lemma 5 and Lemma 6, we can establish the correctness and complexity of Algorithm 3. Specifically, the following theorem holds
Theorem 3.
Given a lattice basis $\mathbf{B}=[\mathbf{a}\ \mathbf{b}]\in\mathbb{Z}^{2\times 2}$ with $\mathbf{a}=(a_{1}\ a_{2})^{T}$, $\mathbf{b}=(b_{1}\ b_{2})^{T}$, and #$\#(\mathbf{a},\mathbf{b})=n$, Algorithm 3 outputs a basis $\mathbf{B}^{\prime}=[\mathbf{a}^{\prime}\ \mathbf{b}^{\prime}]$ satisfying $\|\mathbf{a}^{\prime}\|=\lambda_{1}(L(\mathbf{B}))$ and $\|\mathbf{b}^{\prime}\|=\lambda_{2}(L(\mathbf{B}))$ with a time complexity of $O(n^{2})$.
Proof.
If the input lattice basis satisfies $a_{1}a_{2}b_{1}b_{2}>0$, we assume that the reduced lattice basis sequence in the first While loop of Algorithm 3 is as follows:
$\mathbf{B}^{(0)}=[\mathbf{a}^{(0)}\ \mathbf{b}^{(0)}]=\begin{pmatrix}a_{1}^{(0)}&b_{1}^{(0)}\\ a_{2}^{(0)}&b_{2}^{(0)}\end{pmatrix}=\begin{pmatrix}a_{1}&b_{1}\\ a_{2}&b_{2}\end{pmatrix}\to\mathbf{B}^{(1)}=[\mathbf{a}^{(1)}\ \mathbf{b}^{(1)}]=\begin{pmatrix}a_{1}^{(1)}&b_{1}^{(1)}\\ a_{2}^{(1)}&b_{2}^{(1)}\end{pmatrix}\to\cdots$
$\to\mathbf{B}^{(k-1)}=[\mathbf{a}^{(k-1)}\ \mathbf{b}^{(k-1)}]=\begin{pmatrix}a_{1}^{(k-1)}&b_{1}^{(k-1)}\\ a_{2}^{(k-1)}&b_{2}^{(k-1)}\end{pmatrix}\to\mathbf{B}^{(k)}=[\mathbf{a}^{(k)}\ \mathbf{b}^{(k)}]=\begin{pmatrix}a_{1}^{(k)}&b_{1}^{(k)}\\ a_{2}^{(k)}&b_{2}^{(k)}\end{pmatrix},$ (12)
where $\mathbf{a}^{(i)}=\mathbf{b}^{(i-1)}$ for $i=1,\cdots,k$, and, with loss of generality, we assume $k$ is even. By Lemma 5, we have $a_{1}^{(k)}a_{2}^{(k)}b_{1}^{(k)}b_{2}^{(k)}\leq 0$ and
$\|\mathbf{a}^{(k)}\|=\|\mathbf{b}^{(k-1)}\|<\frac{1}{2}\|\mathbf{a}^{(k-2)}\|<\cdots<\frac{1}{2^{k/2}}\|\mathbf{a}\|,$
$\|\mathbf{a}^{(k-1)}\|<\frac{1}{2}\|\mathbf{a}^{(k-3)}\|<\cdots<\frac{1}{2^{(k-2)/2}}\|\mathbf{a}^{(1)}\|=\frac{1}{2^{(k-2)/2}}\|\mathbf{b}\|$
$\|\mathbf{b}^{(k)}\|<\|\mathbf{a}^{(k-1)}\|\cdot\|\mathbf{b}^{(k-1)}\|<\frac{1}{2^{k/2}}\|\mathbf{a}\|\cdot\frac{1}{2^{(k-2)/2}}\|\mathbf{b}\|=\frac{1}{2^{k-1}}\|\mathbf{a}\|\cdot\|\mathbf{b}\|.$
Since $1\leq\min\{\|\mathbf{a}^{(k)}\|,\|\mathbf{b}^{(k)}\|\}<\frac{1}{2^{k-1}}|\mathbf{a}\|\cdot\|\mathbf{b}\|$, it follows that $k<\log\|\mathbf{a}\|+\log\|\mathbf{b}\|+1$ is finite.
Further, if $(|a_{1}^{(k)}|-|a_{2}^{(k)}|)(|b_{1}^{(k)}|-|b_{2}^{(k)}|)>0$, then during the second While loop, we assume the reduced lattice basis sequence is
$\mathbf{B}^{(k)}=[\mathbf{a}^{(k)}\ \mathbf{b}^{(k)}]=\begin{pmatrix}a_{1}^{(k)}&b_{1}^{(k)}\\ a_{2}^{(k)}&b_{2}^{(k)}\end{pmatrix}\to\mathbf{B}^{(k+1)}=[\mathbf{a}^{(k+1)}\ \mathbf{b}^{(k+1)}]=\begin{pmatrix}a_{1}^{(k+1)}&b_{1}^{(k+1)}\\ a_{2}^{(k+1)}&b_{2}^{(k+1)}\end{pmatrix}\to\cdots$
$\to\mathbf{B}^{(k+s-1)}=[\mathbf{a}^{(k+s-1)}\ \mathbf{b}^{(k+s-1)}]=\begin{pmatrix}a_{1}^{(k+s-1)}&b_{1}^{(k+s-1)}\\ a_{2}^{(k+s-1)}&b_{2}^{(k+s-1)}\end{pmatrix}\to\mathbf{B}^{(k+s)}=[\mathbf{a}^{(k+s)}\ \mathbf{b}^{(k+s)}]=\begin{pmatrix}a_{1}^{(k+s)}&b_{1}^{(k+s)}\\ a_{2}^{(k+s)}&b_{2}^{(k+s)}\end{pmatrix},$ (13)
where $\mathbf{a}^{(k+i)}=\mathbf{b}^{(k+i-1)}$ for $i=1,\cdots,s$, and, without loss of generality, we assume $s$ is even. Then, by Lemma 6, we have $a_{1}^{(k+i)}a_{2}^{(k+i)}b_{1}^{(k+i)}b_{2}^{(k+i)}\leq 0$ for $i=0,\cdots,s$,
$1\leq|a_{1}^{(k+s)}|-|a_{2}^{(k+s)}|=|b_{1}^{(k+(s-1))}|-|b_{2}^{(k+(s-1))}|<\frac{|a_{1}^{(k+(s-2))}|-|a_{2}^{(k+(s-2))}|}{2}<\cdots<\frac{|a_{1}^{(k)}|-|a_{2}^{(k)}|}{2^{s/2}},$
and $|b_{1}^{(k+s)}|-|b_{2}^{(k+s)}|\leq 0.$ Consequently, after the second While loop, we have
$a_{1}^{(k+s)}a_{2}^{(k+s)}b_{1}^{(k+s)}b_{2}^{(k+s)}\leq 0\wedge(|a_{1}^{(k+s)}|-|a_{2}^{(k+s)}|)(|b_{1}^{(k+s)}|-|b_{2}^{(k+s)}|)\leq 0,$
and thus, the lattice basis $\mathbf{B}^{(k+s)}=[\mathbf{a}^{(k+s)}\ \mathbf{b}^{(k+s)}]$ is reduced. Finally, by Theorem 1, the returned basis $\mathbf{B}^{\prime}=[\mathbf{a}^{\prime}\ \mathbf{b}^{\prime}]$ of Algorithm 3 satisfies $\|\mathbf{a}^{\prime}\|=\lambda_{1}(L(\mathbf{B}))$ and $\|\mathbf{b}^{\prime}\|=\lambda_{2}(L(\mathbf{B}))$
Now, we estimate the time complexity. Without loss of generality, we assume $\|\mathbf{a}^{(0)}\|\geq\|\mathbf{b}^{(0)}\|$, #$\#\mathbf{a}^{(i)}=n_{i}$ and #$\#q^{(i)}=\ell_{i}$ for $i=1,\cdots,k,k+1,\cdots,k+s$. For the first While loop, since $\mathbf{a}^{(i-1)}=q^{(i)}\mathbf{b}^{(i-1)}+\mathbf{b}^{(i)}=q^{(i)}\mathbf{a}^{(i)}+\mathbf{a}^{(i+1)}$, where
$q^{(i)}\in\left\{\left\lfloor\frac{a_{1}^{(i-1)}}{b_{1}^{(i-1)}}\right\rfloor_{0}=\left\lfloor\frac{a_{1}^{(i-1)}}{a_{1}^{(i)}}\right\rfloor_{0},\left\lfloor\frac{a_{1}^{(i-1)}}{b_{1}^{(i-1)}}\right\rfloor_{0}+1=\left\lfloor\frac{a_{1}^{(i-1)}}{a_{1}^{(i)}}\right\rfloor_{0}+1,\right.$
$\left.\left\lfloor\frac{a_{2}^{(i-1)}}{b_{2}^{(i-1)}}\right\rfloor_{0}=\left\lfloor\frac{a_{2}^{(i-1)}}{a_{2}^{(i)}}\right\rfloor_{0},\left\lfloor\frac{a_{2}^{(i-1)}}{b_{2}^{(i-1)}}\right\rfloor_{0}+1=\left\lfloor\frac{a_{2}^{(i-1)}}{a_{2}^{(i)}}\right\rfloor_{0}+1\right\},$ (14)
we have $n_{i-1}=n_{i}+\ell_{i}$. Therefore, in each loop, the time complexity of division is bounded by $O(n_{i-1}\ell_{i})$ and the time complexity of multiplication is bounded by $O(n_{i}\ell_{i})$. Consequently, the total complexity of the first While loop is $T_{1}=O(\sum_{i=1}^{k}(n_{i-1}\ell_{i}+n_{i}\ell_{i}))=O(\sum_{i=1}^{k}(n_{i}+n_{i-1})(n_{i-1}-n_{i}))=O(\sum_{i=1}^{k}(n_{i-1}^{2}-n_{i}^{2}))=O(n_{0}^{2}-n_{k}^{2})$. Similarly, for the second While loop, the total complexity is $T_{2}=O(\sum_{i=k+1}^{k+s}(n_{i-1}\ell_{i}+n_{i}\ell_{i}))=O(\sum_{i=k+1}^{k+s}(n_{i}+n_{i-1})(n_{i-1}-n_{i}))=O(\sum_{i=k+1}^{k+s}(n_{i-1}^{2}-n_{i}^{2}))=O(n_{k}^{2}-n_{k+s}^{2})$. Therefore, the total complexity of Algorithm 3 is $T=O(T_{1}+T_{2})=O(n_{0}^{2})=O(n^{2})$. ∎
Remark 1.
According to the property established in Theorem 2, the $\ell_{2}$-shortest basis $[\mathbf{u}\ \mathbf{v}]$ can be obtained via a minor modification of Algorithm 3. Specifically, by replacing Steps 1114 with the following:
$\mathbf{u}=\arg\min\left\{\|\mathbf{a}\|_{2},\ \|\mathbf{b}\|_{2},\ \|\mathbf{a}+\mathbf{b}\|_{2},\ \|\mathbf{a}-\mathbf{b}\|_{2}\right\},$
and setting
$\mathbf{v}=\mathbf{x}-q\mathbf{u},\quad\text{where }\mathbf{x}\in\{\mathbf{a},\mathbf{b}\}\setminus\{\mathbf{u}\}\text{ and }q=\left\lceil\frac{\langle\mathbf{x},\mathbf{u}\rangle}{\langle\mathbf{u},\mathbf{u}\rangle}\right\rfloor.$
This substitution yields a basis satisfying the $\ell_{2}$-shortest condition. The same remark is also applicable to the upcoming Algorithm 5 (HVecSBP).
V Optimized algorithm and its analysis
For two integers with bit lengths of $n$, the well-known Half-GCD algorithm [10, 11] can efficiently find two integers of approximately $n/2$-bit length that share the same common divisor. Inspired by the Half-GCD algorithm, this section first explores the HVec algorithm, which rapidly reduces a lattice basis to a new basis with bit lengths approximately halved from the original, along with a complexity analysis. Furthermore, we leverage this algorithm to optimize and accelerate the CrossEuc algorithm presented in the previous section.
V-A HVec algorithm and its analysis
Analogous to the case of two integers, our design for the input of two vectors is based on the following key observation:
Assume $[\mathbf{a}^{\prime}\ \mathbf{b}^{\prime}]$ and $[\mathbf{c}^{\prime}\ \mathbf{d}^{\prime}]$ are two different bases for the same lattice, i.e.,
$[\mathbf{a}^{\prime}\ \mathbf{b}^{\prime}]=\begin{pmatrix}a_{1}^{\prime}&b_{1}^{\prime}\\ a_{2}^{\prime}&b_{2}^{\prime}\end{pmatrix}=\begin{pmatrix}c_{1}^{\prime}&d_{1}^{\prime}\\ c_{2}^{\prime}&d_{2}^{\prime}\end{pmatrix}\mathbf{M}^{\prime}=[\mathbf{c}^{\prime}\ \mathbf{d}^{\prime}]\mathbf{M}^{\prime}$ (15)
for some unimodular matrix $\mathbf{M}^{\prime}$. The two new vectors
$[\mathbf{a}\ \mathbf{b}]=\begin{pmatrix}a_{1}&b_{1}\\ a_{2}&b_{2}\end{pmatrix}=2^{n_{\ell}}\begin{pmatrix}a_{1}^{\prime}&b_{1}^{\prime}\\ a_{2}^{\prime}&b_{2}^{\prime}\end{pmatrix}+\begin{pmatrix}a_{1}^{\prime\prime}&b_{1}^{\prime\prime}\\ a_{2}^{\prime\prime}&b_{2}^{\prime\prime}\end{pmatrix}=2^{n_{\ell}}[\mathbf{a}^{\prime}\ \mathbf{b}^{\prime}]+[\mathbf{a}^{\prime\prime}\ \mathbf{b}^{\prime\prime}],$ (16)
where $[\mathbf{a}^{\prime\prime}\ \mathbf{b}^{\prime\prime}]$ represents the $n_{\ell}$ least significant bits, and $[\mathbf{a}^{\prime}\ \mathbf{b}^{\prime}]$ refers to the $n_{h}=\#(\mathbf{a},\mathbf{b})-n_{\ell}$ most significant bits of $[\mathbf{a}\ \mathbf{b}]$. Additionally, the vectors $[\mathbf{c}\ \mathbf{d}]$ satisfy
$[\mathbf{a}\ \mathbf{b}]=\begin{pmatrix}a_{1}&b_{1}\\ a_{2}&b_{2}\end{pmatrix}=\begin{pmatrix}c_{1}&d_{1}\\ c_{2}&d_{2}\end{pmatrix}\mathbf{M}^{\prime}=[\mathbf{c}\ \mathbf{d}]\mathbf{M}^{\prime}$ (17)
for the same unimodular matrix $\mathbf{M}^{\prime}$. Then, we have
$[\mathbf{c}\ \mathbf{d}]=[\mathbf{a}\ \mathbf{b}](\mathbf{M}^{\prime})^{-1}=2^{n_{\ell}}[\mathbf{c}^{\prime}\ \mathbf{d}^{\prime}]+[\mathbf{a}^{\prime\prime}\ \mathbf{b}^{\prime\prime}](\mathbf{M}^{\prime})^{-1}$ (18)
Building on the above fact, we can halve the bit-length of $[\mathbf{a}\ \mathbf{b}]$ by adopting a recursive reduction strategy. Algorithm 4 provides the details of our design. We now provide a rigorous theoretical analysis to establish the correctness and complexity of Algorithm 4. That is, we argue the following result:
Algorithm 4 HVec($\mathbf{a}\ \mathbf{b}$)
0:  Two vectors $\mathbf{a}=(a_{1},a_{2})^{T},\mathbf{b}=(b_{1},b_{2})^{T}$ and #$\#(\mathbf{a},\mathbf{b})=n,\underline{\#}(\mathbf{a},\mathbf{b})>\left\lfloor\frac{n}{2}\right\rfloor+1,\underline{\#}(\mathbf{a}+\mathbf{b},\mathbf{a}-\mathbf{b})>\lfloor\frac{n}{2}\rfloor+1$
0:  Two vectors $\mathbf{c}=(c_{1},c_{2})^{T},\mathbf{d}=(d_{1},d_{2})^{T}$ and an unimodular matrix $\mathbf{M}$ such that $[\mathbf{a}\ \mathbf{b}]=[\mathbf{c}\ \mathbf{d}]\mathbf{M}$. Meanwhile, either (#$(\#(\mathbf{c},\mathbf{d})\leq n)\wedge(\underline{\#}(\mathbf{c},\mathbf{d})>\lfloor\frac{n}{2}\rfloor+1)\wedge(\underline{\#}(\mathbf{c}+\mathbf{d},\mathbf{c}-\mathbf{d})\leq\lfloor\frac{n}{2}\rfloor+1)$, or $[\mathbf{c}\ \mathbf{d}]$ is reduced.
1:  $\mathbf{c}:=\mathbf{a},\mathbf{d}:=\mathbf{b}$, $\mathbf{M}:=\mathbf{I}$
2:  if $c_{1}c_{2}d_{1}d_{2}\leq 0\wedge(|c_{1}|-|c_{2}|)(|d_{1}|-|d_{2}|)\leq 0$ then
3:     return  $\mathbf{c},\mathbf{d},\mathbf{M}$
4:  end if
5:  $n:=\#(\mathbf{c},\mathbf{d})$
6:  $s:=\lfloor\frac{n}{2}\rfloor+1$
7:  if #¯$\underline{\#}(\mathbf{c},\mathbf{d})\geq\lfloor\frac{3}{4}n\rfloor+2$, then
8:   $n_{\ell 1}:=\lfloor n/2\rfloor$
9:   $a_{1}^{\prime}:=\lfloor c_{1}/2^{n_{\ell 1}}\rfloor$, $a_{1}^{\prime\prime}:=c_{1}-2^{n_{\ell 1}}a_{1}^{\prime}$, $a_{2}^{\prime}:=\lfloor c_{2}/2^{n_{\ell 1}}\rfloor$, $a_{2}^{\prime\prime}:=c_{2}-2^{n_{\ell 1}}a_{2}^{\prime}$
10:   $b_{1}^{\prime}:=\lfloor d_{1}/2^{n_{\ell 1}}\rfloor$, $b_{1}^{\prime\prime}:=d_{1}-2^{n_{\ell 1}}b_{1}^{\prime}$, $b_{2}^{\prime}:=\lfloor d_{2}/2^{n_{\ell 1}}\rfloor$, $b_{2}^{\prime\prime}:=d_{2}-2^{n_{\ell 1}}b_{2}^{\prime}$
11:   $\mathbf{a}^{\prime}:=(a_{1}^{\prime},a_{2}^{\prime})^{T},\mathbf{b}^{\prime}:=(b_{1}^{\prime},b_{2}^{\prime})^{T}$, $\mathbf{a}^{\prime\prime}:=(a_{1}^{\prime\prime},a_{2}^{\prime\prime})^{T},\mathbf{b}^{\prime\prime}:=(b_{1}^{\prime\prime},b_{2}^{\prime\prime})^{T}$
12:   $(\mathbf{c}^{\prime},\mathbf{d}^{\prime},\mathbf{M}_{1}^{\prime})\leftarrow\textbf{HVec}(\mathbf{a}^{\prime},\mathbf{b}^{\prime})$
13:   $[\mathbf{c}\ \mathbf{d}]:=2^{n_{\ell 1}}[\mathbf{c}^{\prime}\ \mathbf{d}^{\prime}]+[\mathbf{a}^{\prime\prime}\ \mathbf{b}^{\prime\prime}](\mathbf{M}_{1}^{\prime})^{-1}$
14:   $\mathbf{M}:=\mathbf{M}_{1}^{\prime}\mathbf{M}$
15:  end if
16:  while #$\#(\mathbf{c},\mathbf{d})>\lfloor 3n/4\rfloor+1$ and #¯$\underline{\#}(\mathbf{c}-\mathbf{d},\mathbf{c}+\mathbf{d})>s$ do
17:   if ($c_{1}c_{2}d_{1}d_{2}\leq 0\wedge(|c_{1}|-|c_{2}|)(|d_{1}|-|d_{2}|)\leq 0$), return $\mathbf{c},\mathbf{d},\mathbf{M}$
18:   if ($c_{1}c_{2}d_{1}d_{2}>0$),
19:    $(\mathbf{c},\mathbf{d},q)\leftarrow\textbf{UMTrans1}(\mathbf{c},\mathbf{d})$
20:   else
21:    $(\mathbf{c},\mathbf{d},q)\leftarrow\textbf{UMTrans2}(\mathbf{c},\mathbf{d})$
22:   if (#$\#\mathbf{d}\leq s$), $q:=q-1,\mathbf{d}:=\mathbf{c}+\mathbf{d}$
23:   $\mathbf{M}:=\begin{pmatrix}q&1\\ 1&0\end{pmatrix}\mathbf{M}$
24:  end while
25:  if #¯$\underline{\#}(\mathbf{c},\mathbf{d})>s+2$, then
26:   $n^{\prime}:=\#(\mathbf{c},\mathbf{d})$, $n_{\ell 2}:=2s-n^{\prime}+1$
27:   $a_{1}^{\prime}:=\lfloor c_{1}/2^{n_{\ell 2}}\rfloor$, $a_{1}^{\prime\prime}:=c_{1}-2^{n_{\ell 2}}a_{1}^{\prime}$, $a_{2}^{\prime}:=\lfloor c_{2}/2^{n_{\ell 2}}\rfloor$, $a_{2}^{\prime\prime}:=c_{2}-2^{n_{\ell 2}}a_{2}^{\prime}$
28:   $b_{1}^{\prime}:=\lfloor d_{1}/2^{n_{\ell 2}}\rfloor$, $b_{1}^{\prime\prime}:=d_{1}-2^{n_{\ell 2}}b_{1}^{\prime}$, $b_{2}^{\prime}:=\lfloor d_{2}/2^{n_{\ell 2}}\rfloor$, $b_{2}^{\prime\prime}:=d_{2}-2^{n_{\ell 2}}b_{2}^{\prime}$
29:   $\mathbf{a}^{\prime}:=(a_{1}^{\prime},a_{2}^{\prime})^{T},\mathbf{b}^{\prime}:=(b_{1}^{\prime},b_{2}^{\prime})^{T}$, $\mathbf{a}^{\prime\prime}:=(a_{1}^{\prime\prime},a_{2}^{\prime\prime})^{T},\mathbf{b}^{\prime\prime}:=(b_{1}^{\prime\prime},b_{2}^{\prime\prime})^{T}$
30:   $(\mathbf{c}^{\prime},\mathbf{d}^{\prime},\mathbf{M}_{2}^{\prime})\leftarrow\textbf{HVec}(\mathbf{a}^{\prime},\mathbf{b}^{\prime})$
31:   $[\mathbf{c}\ \mathbf{d}]:=2^{n_{\ell 2}}[\mathbf{c}^{\prime}\ \mathbf{d}^{\prime}]+[\mathbf{a}^{\prime\prime}\ \mathbf{b}^{\prime\prime}](\mathbf{M}_{2}^{\prime})^{-1}$
32:   $\mathbf{M}:=\mathbf{M}_{2}^{\prime}\mathbf{M}$
33:  end if
34:  while #¯$\underline{\#}(\mathbf{c}-\mathbf{d},\mathbf{c}+\mathbf{d})>s$ do
35:   if ($c_{1}c_{2}d_{1}d_{2}\leq 0\wedge(|c_{1}|-|c_{2}|)(|d_{1}|-|d_{2}|)\leq 0$), return $\mathbf{a},\mathbf{b},\mathbf{M}$
36:   if ($c_{1}c_{2}d_{1}d_{2}>0$),
37:    $(\mathbf{c},\mathbf{d},q)\leftarrow\textbf{UMTrans1}(\mathbf{c},\mathbf{d})$
38:   else
39:    $(\mathbf{c},\mathbf{d},q)\leftarrow\textbf{UMTrans2}(\mathbf{c},\mathbf{d})$
40:   if (#$\#\mathbf{d}\leq s$), $q:=q-1,\mathbf{d}:=\mathbf{c}+\mathbf{d}$
41:   $\mathbf{M}:=\begin{pmatrix}q&1\\ 1&0\end{pmatrix}\mathbf{M}$
42:  end while
43:  return $\mathbf{c},\mathbf{d},\mathbf{M}$
Theorem 4.
For any input basis $\mathbf{B}=[\mathbf{a}\ \mathbf{b}]\in\mathbb{Z}^{2\times 2}$ with #$\#(\mathbf{a},\mathbf{b})=n$, Algorithm 4 will output a new basis $\mathbf{B}^{\prime}=[\mathbf{c}\ \mathbf{d}]$ and an unimodular matrix $\mathbf{M}$ such that $[\mathbf{a}\ \mathbf{b}]=[\mathbf{c}\ \mathbf{d}]\mathbf{M}$ in time $T(n)=O\left(M(n)\log n\right)$, where $M(n)$ refers to the time complexity of multiplying two $n$-bit integers. Additionally, the basis $[\mathbf{c}\ \mathbf{d}]$ satisfies one of the following two conditions:
  1. 1.
    $[\mathbf{c},\mathbf{d}]$ is reduced, or
  2. 2.
    #¯$\underline{\#}(\mathbf{c},\mathbf{d})>\lfloor\frac{n}{2}\rfloor+1$, and #¯$\underline{\#}(\mathbf{c}-\mathbf{d},\mathbf{c}+\mathbf{d})\leq\lfloor\frac{n}{2}\rfloor+1$.
Proof.
First, we argue that the algorithm will inevitably terminate. Assume
$(\mathbf{c}^{new},\ \mathbf{d}^{new},q)\leftarrow\textbf{UMTrans1}(\mathbf{c}\ \mathbf{d})(\textrm{resp.}\ \textbf{UMTrans2}(\mathbf{c}\ \mathbf{d})).$
Then, by Lemma 5 and Lemma 6, if $[\mathbf{c}^{new},\ \mathbf{d}^{new}]$ is not reduced, the following inequalities hold except for a special case where $c_{1}c_{2}d_{1}d_{2}>0$ and $c_{1}^{new}c_{2}^{new}d_{1}^{new}d_{2}^{new}\leq 0$, which occurs only once during execution and can be ignored:
$(\mathbf{c}^{new}=\mathbf{d})\wedge(\|\mathbf{d}^{new}\|\leq\min\{\|\mathbf{c}-\mathbf{d}\|,\|\mathbf{c}+\mathbf{d}\|\})\wedge(\|\mathbf{d}^{new}\|<\|\mathbf{c}^{new}\|)\wedge(\|\mathbf{d}^{new}\|<\frac{1}{2}\|\mathbf{c}\|),$
which implies that
#$\#(\mathbf{c}^{(new)},\mathbf{d}^{(new)})<\#(\mathbf{c},\mathbf{d}),\ \#\mathbf{d}^{(new)}\leq\underline{\#}(\mathbf{c}+\mathbf{d},\mathbf{c}-\mathbf{d}),\ \#\mathbf{d}^{(new)}<\#\mathbf{c}-1.$ (19)
Meanwhile, steps 22 and 40 ensure that #¯$\underline{\#}(\mathbf{c}^{(new)},\mathbf{d}^{(new)})>s=\lfloor\frac{n}{2}\rfloor+1$ Consequently, the While conditions in Step 16 and Step 34 will eventually be violated. That is, Algorithm 4 will terminate. Additionally, by equations (15)-(18) and the properties of the unimodular transformations UMTrans1 and UMTrans2, the output $(\mathbf{c},\mathbf{d},\mathbf{M})$ satisfies $[\mathbf{a}\ \mathbf{b}]=[\mathbf{c}\ \mathbf{d}]\mathbf{M}$ and one of the following conditions holds:
(1) $[\mathbf{c}\ \mathbf{d}]$ is reduced, or
(2) #¯$\underline{\#}(\mathbf{c},\mathbf{d})>\lfloor\frac{n}{2}\rfloor+1$ and #¯$\underline{\#}(\mathbf{c}-\mathbf{d},\mathbf{c}+\mathbf{d})\leq\lfloor\frac{n}{2}\rfloor+1$.
Now, we estimate the time complexity. We need to estimate the bit size of the numbers and the time complexity of each while loop during execution. According to our recursive design, the underlying operations are the unimodular transformations UMTrans1 and UMTrans2. For each recursive invocation $(\mathbf{c}^{\prime},\mathbf{d}^{\prime},\mathbf{M}^{\prime})\leftarrow\textbf{HVec}(\mathbf{a}^{\prime}\ \mathbf{b}^{\prime})$, we adopt the same notations as in the proof of Theorem 3 for brevity. Assume that $\|\mathbf{a}^{\prime}\|\geq\|\mathbf{b}^{\prime}\|$ and the lattice basis sequence is given by
$\mathbf{B}^{(0)}=[\mathbf{a}^{(0)}\ \mathbf{b}^{(0)}]=[\mathbf{a}^{\prime}\ \mathbf{b}^{\prime}]\to\mathbf{B}^{(1)}=[\mathbf{a}^{(1)}\ \mathbf{b}^{(1)}]\to\cdots\to\mathbf{B}^{(m)}=[\mathbf{a}^{(m)}\ \mathbf{b}^{(m)}]=[\mathbf{c}^{\prime}\ \mathbf{d}^{\prime}],$
where, for $i=1,\cdots,m$,
$\mathbf{a}^{(i)}=\mathbf{b}^{(i-1)},\mathbf{a}^{(i-1)}=q^{(i)}\mathbf{b}^{(i-1)}+\mathbf{b}^{(i)}=q^{(i)}\mathbf{a}^{(i)}+\mathbf{a}^{(i+1)}$ (20)
and $q^{(i)}$ is defined in equation (IV-B). Then $[\mathbf{a}^{\prime}\ \mathbf{b}^{\prime}]=[\mathbf{c}^{\prime}\ \mathbf{d}^{\prime}]\mathbf{M}^{\prime}$ with
$\mathbf{M}^{\prime}=\begin{pmatrix}q^{(m)}&1\\ 1&0\end{pmatrix}\begin{pmatrix}q^{(m-1)}&1\\ 1&0\end{pmatrix}\cdot\cdots\cdot\begin{pmatrix}q^{(1)}&1\\ 1&0\end{pmatrix}.$ (21)
Furthermore, by Lemma 5 and Lemma 6, the following inequalities hold for almost all $i\in\{1,\dots,m\}$, with at most one exception, which does not affect the proof and can be ignored:
$\|\mathbf{a}^{(i+1)}\|=\|\mathbf{b}^{(i)}\|<\frac{1}{2}\|\mathbf{a}^{(i-1)}\|\quad\text{and}\quad\|\mathbf{b}^{(i)}\|<\|\mathbf{a}^{(i)}\|.$
Then, the equation (20) indicate that #$\#\mathbf{a}^{(i-1)}=\#q^{(i)}+\#\mathbf{a}^{(i)}$. While, the equation (21) implies that #$\#\mathbf{M}^{\prime}<\sum_{i=1}^{m}\#q^{(i)}+1$. Hence,
#$\#\mathbf{a}^{(0)}=\#q^{(1)}+\#\mathbf{a}^{(1)}=\sum_{i=1}^{m}\#q^{(i)}+\#\mathbf{a}^{(i)}>\#\mathbf{M}^{\prime}-1+\#\mathbf{c}^{\prime}$
#$\implies\#\mathbf{M}^{\prime}<\#\mathbf{a}^{(0)}-\#\mathbf{c}^{\prime}+1\leq\#(\mathbf{a}^{\prime},\mathbf{b}^{\prime})-\underline{\#}(\mathbf{c}^{\prime},\mathbf{d}^{\prime})+1.$ (22)
Based on the above observation, in step 7-step 15, we have
#$\#(\mathbf{c}^{\prime},\mathbf{d}^{\prime})\leq n-n_{\ell 1}\wedge\underline{\#}(\mathbf{c}^{\prime},\mathbf{d}^{\prime})>\lfloor\frac{n-n_{\ell 1}}{2}\rfloor+1\wedge\underline{\#}(\mathbf{c}^{\prime}+\mathbf{d}^{\prime},\mathbf{c}^{\prime}-\mathbf{d}^{\prime})\leq\lfloor\frac{n-n_{\ell 1}}{2}\rfloor+1\wedge$
#$\#\mathbf{M}_{1}^{\prime}\leq\#(\mathbf{a}^{\prime},\mathbf{b}^{\prime})-\underline{\#}(\mathbf{c}^{\prime},\mathbf{d}^{\prime})+1<n-n_{\ell 1}-(\lfloor\frac{n-n_{\ell 1}}{2}\rfloor+1)+1.$
Thus, the equation $[\mathbf{c}\ \mathbf{d}]=2^{n_{\ell 1}}[\mathbf{c}^{\prime}\ \mathbf{d}^{\prime}]+[\mathbf{a}^{\prime\prime}\ \mathbf{b}^{\prime\prime}](\mathbf{M}_{1}^{\prime})^{-1}$ in step 13 implies that
#$\#(\mathbf{c},\mathbf{d})\leq\max\{n_{\ell 1}+\#(\mathbf{c}^{\prime},\mathbf{d}^{\prime}),n_{\ell 1}+\#\mathbf{M}_{1}^{\prime}\}+1=n+1,$
#¯$\underline{\#}(\mathbf{c}\pm\mathbf{d})=\underline{\#}(2^{n_{\ell 1}}(\mathbf{c}^{\prime}+\mathbf{d}^{\prime})+(\mathbf{a}^{\prime\prime}+\mathbf{b}^{\prime\prime})(\mathbf{M}_{1}^{\prime})^{-1},2^{n_{\ell 1}}(\mathbf{c}^{\prime}-\mathbf{d}^{\prime})+(\mathbf{a}^{\prime\prime}-\mathbf{b}^{\prime\prime})(\mathbf{M}_{1}^{\prime})^{-1})$
$=\min\{\#(2^{n_{\ell 1}}(\mathbf{c}^{\prime}+\mathbf{d}^{\prime})+(\mathbf{a}^{\prime\prime}+\mathbf{b}^{\prime\prime})(\mathbf{M}_{1}^{\prime})^{-1}),\#(2^{n_{\ell 1}}(\mathbf{c}^{\prime}-\mathbf{d}^{\prime})+(\mathbf{a}^{\prime\prime}-\mathbf{b}^{\prime\prime})(\mathbf{M}_{1}^{\prime})^{-1}))\}$
$=\min\{\max\{n_{\ell 1}+\#(\mathbf{c}^{\prime}+\mathbf{d}^{\prime}),\#(\mathbf{a}^{\prime\prime}+\mathbf{b}^{\prime\prime})+\#\mathbf{M}_{1}^{\prime}\}+1,\max\{n_{\ell 1}+\#(\mathbf{c}^{\prime}-\mathbf{d}^{\prime}),\#(\mathbf{a}^{\prime\prime}-\mathbf{b}^{\prime\prime})+\#\mathbf{M}_{1}^{\prime}\}+1\}$
$\leq\max\{n_{\ell 1}+\underline{\#}(\mathbf{c}^{\prime}\pm\mathbf{d}^{\prime}),\max\{\#\mathbf{a}^{\prime\prime},\#\mathbf{b}^{\prime\prime}\}+1+\#\mathbf{M}_{1}^{\prime}\}+1$
$\leq\max\{n_{\ell 1}+\lfloor\frac{n-n_{\ell 1}}{2}\rfloor+1,n_{\ell 1}+1+n-n_{\ell 1}-(\lfloor\frac{n-n_{\ell 1}}{2}\rfloor+1)+1\}$
$=\max\{n_{\ell 1}+\lfloor\frac{n_{\ell 1}}{2}\rfloor+1,n-\lfloor\frac{n_{\ell 1}}{2}\rfloor+1\}<\frac{3n}{4}+3<\lfloor\frac{3n}{4}\rfloor+4.$
By the equation (19), the bit length of the new vector generated by the unimodular matrix transformation UMTrans1 (resp. UMTrans2) is less that $\min\{\underline{\#}(\mathbf{c}+\mathbf{d},\mathbf{c}-\mathbf{d}),\#\mathbf{c}-1\}$. Therefore, the while loop in step 16 will terminate after at most $8$ iterations and thus the time complexity is $O(M(n))$. Similarly, in step 25-step 33, we have
#$\#\mathbf{M}_{2}^{\prime}\leq\#(\mathbf{a}^{\prime},\mathbf{b}^{\prime})-\underline{\#}(\mathbf{c}^{\prime},\mathbf{d}^{\prime})+1<n^{\prime}-n_{\ell 2}-(\lfloor\frac{n^{\prime}-n_{\ell 2}}{2}\rfloor+1)+1=n^{\prime}-s,$
#$\#(\mathbf{c},\mathbf{d})\leq\max\{n_{\ell 2}+\#(\mathbf{c}^{\prime},\mathbf{d}^{\prime}),n_{\ell 2}+\#\mathbf{M}_{2}^{\prime}\}+1=n^{\prime}+1,$
#¯$\underline{\#}(\mathbf{c}\pm\mathbf{d})=\underline{\#}(2^{n_{\ell 2}}(\mathbf{c}^{\prime}+\mathbf{d}^{\prime})+(\mathbf{a}^{\prime\prime}+\mathbf{b}^{\prime\prime})(\mathbf{M}_{2}^{\prime})^{-1},2^{n_{\ell 2}}(\mathbf{c}^{\prime}-\mathbf{d}^{\prime})+(\mathbf{a}^{\prime\prime}-\mathbf{b}^{\prime\prime})(\mathbf{M}_{2}^{\prime})^{-1})<s+2.$
By the equation (19), the bit length of the new vector generated by the unimodular matrix transformation UMTrans1 (resp. UMTrans2) is less that $\min\{\underline{\#}(\mathbf{c}+\mathbf{d},\mathbf{c}-\mathbf{d}),\#\mathbf{c}-1\}$. Therefore, the while loop in step 34 will terminate after at most $4$ iterations and thus the time complexity is $O(M(n))$. Overall, the total time complexity is
$T(n)=2T\left(\frac{n}{2}\right)+O(M(n))=\left(2^{2}T\left(\frac{n}{2^{2}}\right)+O\left(2\cdot M\left(\frac{n}{2}\right)\right)\right)+O(M(n))=O\left(\sum_{i=0}^{\log n}2^{i}M\left(\frac{n}{2^{i}}\right)\right).$
Since $M\left(\frac{n}{2^{i}}\right)\leq\frac{1}{2^{i}}M\left(n\right)$ for $i=0,\cdots,\log n$,
$T(n)=O\left(\sum_{i=0}^{\log n}2^{i}M\left(\frac{n}{2^{i}}\right)\right)=O\left(M(n)\sum_{i=0}^{\log n}1\right)=O(M(n)\log n).$
V-B HVecSBP algorithm and its complexity analysis
By cyclically invoking the proposed HVec algorithm, the reduced basis can ultimately be obtained. This design, referred to as HVecSBP, is detailed in Algorithm 5. Clearly, by Theorem 4, the HVecSBP algorithm will output a reduced basis with a worst-case time complexity of
$O\left(T(n)+T\left(\frac{n}{2}\right)+\cdots+T\left(\frac{n}{2^{\log n}}\right)\right)=O\left(M(n)\log n+M\left(\frac{n}{2}\right)\log\frac{n}{2}+\cdots+M\left(\frac{n}{2^{\log n}}\right)\log\frac{n}{2^{\log n}}\right)$
$=O\left(M(n)\log n+\frac{1}{2}M\left(n\right)\log n+\cdots+\frac{1}{2^{\log n}}M\left(n\right)\log n\right)=O(M(n)\log n)$
Algorithm 5 HVecSBP($\mathbf{a},\mathbf{b}$)
0:  A base $[\mathbf{a}\ \mathbf{b}]$ with $\mathbf{a}=(a_{1}\ a_{2})^{T}$, $\mathbf{b}=(b_{1}\ b_{2})^{T}$
0:  A reduced base $[\mathbf{a}\ \mathbf{b}]$
1:  $\mathbf{M}=\mathbf{I}$
2:  While(1)
3:   if ($a_{1}a_{2}b_{1}b_{2}\leq 0\wedge(|a_{1}|-|a_{2}|)(|b_{1}|-|b_{2}|)\leq 0$) break
4:   if $\|\mathbf{a}\|<\|\mathbf{b}\|$, Swap($\mathbf{a},\mathbf{b}$)
5:   $n:=\#\mathbf{a}$
6:   if #$\#\mathbf{b}\leq\left\lfloor\frac{n}{2}\right\rfloor+1$
7:    if ($|a_{1}|\geq|a_{2}|\wedge b_{1}\neq 0$), $q:=\left\lfloor\frac{a_{1}}{b_{1}}\right\rfloor_{o}$
8:    else $q:=\left\lfloor\frac{a_{2}}{b_{2}}\right\rfloor_{o}$
9:    ($\mathbf{a},\mathbf{b}$) := ($\mathbf{b},\mathbf{a}-q\mathbf{b}$)
10:    continue
11:   $(\mathbf{a},\mathbf{b},\mathbf{M})\leftarrow\textbf{HVec}(\mathbf{a},\mathbf{b})$
12:   if #$\#(\mathbf{a}+\mathbf{b})\leq\lfloor\frac{n}{2}\rfloor+1$, ($\mathbf{a},\mathbf{b}$) := ($\mathbf{b},\mathbf{a}+\mathbf{b}$)
13:   else if #$\#(\mathbf{a}-\mathbf{b})\leq\lfloor\frac{n}{2}\rfloor+1$, ($\mathbf{a},\mathbf{b}$) := ($\mathbf{b},\mathbf{a}-\mathbf{b}$)
14:  If (($a_{1}b_{1}=0\wedge sgn(a_{2})\neq sgn(b_{2}))$ or ($a_{2}b_{2}=0\wedge sgn(a_{1})\neq sng(b_{1})$)), $\mathbf{b}:=-\mathbf{b}$
15:  If $\|\mathbf{a}\|>\|\mathbf{b}\|$, Swap($\mathbf{a},\mathbf{b}$)
16:  If $\mathbf{a}=[0,0]$, Return[$\mathbf{a},\mathbf{b}$]
17:  Else
18:   if ($a_{2}b_{2}\leq 0$)
19:    $\mathbf{b}:=\min\left\{\mathbf{b}-\left\lfloor\frac{|b_{1}|-|b_{2}|}{|a_{1}|+|a_{2}|}\right\rfloor\mathbf{a},\mathbf{b}-\left\lceil\frac{|b_{1}|-|b_{2}|}{|a_{1}|+|a_{2}|}\right\rceil\mathbf{a}\right\}$
20:   else
21:    $\mathbf{b}:=\min\left\{\mathbf{b}-\left\lfloor\frac{|b_{2}|-|b_{1}|}{|a_{1}|+|a_{2}|}\right\rfloor\mathbf{a},\mathbf{b}-\left\lceil\frac{|b_{2}|-|b_{1}|}{|a_{1}|+|a_{2}|}\right\rceil\mathbf{a}\right\}$
22:   Return $[\mathbf{a}\ \mathbf{b}]$
VI Practical Performance Evaluation
VI-A Experimental Methodology
To comprehensively evaluate the practical efficiency of our proposed algorithm CrossEuc and its optimized variant HVecSBP against existing methods, we will conduct extensive experiments considering both metric types and lattice basis forms. In terms of the metric, we compare the practical performance of various algorithms under both the Euclidean and Minkowski metrics. Regarding the input basis, it is evident that the performance of existing algorithms is highly dependent on the structure of the input lattice basis. We focus on two common types: the HNF-form basis $\mathbf{B}=[\mathbf{a}\ \mathbf{b}]=\begin{pmatrix}a&b\\ 0&c\end{pmatrix}$, where $a>b>|c|>0$, and the general-form basis $\mathbf{B}=[\mathbf{a}\ \mathbf{b}]=\begin{pmatrix}a_{1}&b_{1}\\ a_{2}&b_{2}\end{pmatrix}$. For general-form bases, the degree of linear dependence between the two basis vectors clearly affects algorithmic performance. We define the linear dependency measure $\Delta(\mathbf{a},\mathbf{b})$ as:
$\Delta(\mathbf{a},\mathbf{b})=\left|\frac{a_{1}}{b_{1}}-\frac{a_{2}}{b_{2}}\right|.$
It is clear that a smaller $\Delta$ value indicates a higher degree of linear dependency between vectors $\mathbf{a}$ and $\mathbf{b}$.
(1) Evaluation for $\ell_{2}$-shortest basis. We compare five algorithms: our proposed fundamental reduction algorithm CrossEuc, Lagrange reduction algorithm LagRed under the $\ell_{2}$ metric [6, 9] (see Appendix B), Yap’s fundamental algorithm CRS [15] (see Appendix C), our optimized algorithm HVecSBP and Yap’s optimized algorithm (HalfGaussianSBP) [15] (see Appendix D).
  • For HNF-form bases, we assessed the practical performance of the above-mentioned algorithms across different bit sizes of $c$.
  • For general-form bases, we assess the practical performance of the above-mentioned algorithms across different $\Delta$ values.
(2) Evaluation for $\ell_{\infty}$-shortest basis.
  • For HNF-form bases, our algorithm CrossEuc is identical to the continued fraction-based algorithm (ParEuc) [4, 3]. Thus, we only compare three algorithms: the optimized Lagrange algorithm under the $\ell_{\infty}$ metric (GolEuc) [6, 3], the continued fraction-based algorithm (ParEuc) [4, 3], and our optimized algorithm (HVecSBP).
  • For general-form bases, we evaluate the practical performance of the following algorithms across different $\Delta$ values.
    • GolEuc: The optimized Lagrange algorithm under the $\ell_{\infty}$ metric presented in [3].
    • EEA-HNF-ParEuc: An algorithm that first transforms the basis into HNF using the extended Euclidean algorithm (EEA) and then reduces it with the continued fraction-based algorithm (ParEuc). Visually
[…]
    • CrossEuc: Our proposed reduced algorithm without using Hvec .
    • HGCD-HNF-ParEuc: Similar to EEA-HNF-ParEuc, but the basis is transformed into HNF using the HGCD-based extended Euclidean algorithm. Visually
[…]
    • HGCD-HNF-HVecSBP: Similar to HGCD-HNF-ParEuc, but the HNF is reduced using our proposed HVecSBP algorithm. Visually
[…]
    • HVecSBP: Our proposed algorithm applied directly to the input basis without prior transformation to HNF.
VI-B Experimental Environment and Result Analysis
Our experiment are performed on a Ubuntu 22.04 machine with Intel Core i7-9750H, 2.6 GHz CPU and 16 GB RAM and implement the algorithm with C language. The detailed experimental results and their analysis are presented as follows.
VI-B1 Experimental result and analysis for the $\ell_{2}$-shortest basis
  • For the input lattice basis in Hermite Normal Form (HNF), Table I compares the running time of each algorithm as $n_{2}$ varies, where $n_{1}=2\times 10^{5}$ denotes the number of decimal digits in $a$ and $b$, and $n_{2}$ represents the number of decimal digits in $c$. As shown in Table I, smaller values of $n_{2}$ lead to higher computational times for all algorithms. In addition, the following observations can be made:
    • (i)
      For the three fundamental reduction algorithms CrossEuc, CRS, and LagRed, our proposed CrossEuc significantly outperforms the others, with time consumption approximately only 0.19% of CRS and 0.28% of LagRed, achieving speedups of around 500×500$\times$500 × and 350×350$\times$350 ×, respectively.
    • (ii)
      For the two optimized algorithms HalfGaussianSBP and HVecSBP, our optimized algorithm HVecSBP takes only about $7\%$ of the time used by Yap’s HalfGaussianSBP, resulting in a speedup of approximately 14×14$\times$14 ×.
    • (iii)
      For our proposed fundamental reduction algorithm $\mathbf{CrossEuc}$ and its optimized variant $\mathbf{HVecSBP}$, the runtime of HVecSBP constitutes $18.8\%$-$38.1\%$ of CrossEuc’s as $n_{2}$ decreases – achieving a $5.3$ to $2.6$-fold efficiency gain. Remarkably, $\mathbf{CrossEuc}$ itself outperforms Yap’s $\mathbf{HalfGaussianSBP}$, requiring only $35.6\%$-$18.9\%$ of its runtime under diminishing $n_{2}$. This advantage is particularly pronounced for smaller-scale lattice bases (characterized by parameter $n_{1}$).
    TABLE I: Running time of various algorithms for the $\ell_{2}$-shortest basis with varying $n_{2}$ (in seconds)
    length (dec) $n_{1}$  length (dec) $n_{2}$  CrossEuc  CRS  LagRed  HalfGaussianSBP  HVecSBP
    200000 200000 0.000522 0.035561 0.034960 0.072230 0.001405
    200000 180000 0.215646 112.676702 77.524088 0.605316 0.040622
    200000 160000 0.394690 213.510331 146.752125 1.156759 0.078501
    200000 140000 0.557541 296.380581 204.626063 1.689438 0.109761
    200000 120000 0.694540 366.024648 255.231900 2.243587 0.152746
    200000 100000 0.803900 423.430425 296.717711 2.769396 0.197548
    200000 80000 0.902036 464.113459 323.928026 3.277750 0.222285
    200000 60000 0.952363 498.170153 346.013871 3.902740 0.262041
    200000 40000 1.014240 517.602416 358.009027 4.380847 0.312382
    200000 20000 1.037861 529.151738 365.802161 4.944157 0.350908
    200000 10000 1.044446 530.145559 370.650970 5.307488 0.366273
    200000 1000 1.046709 531.093756 371.334628 5.534359 0.394135
    200000 100 1.052189 531.633906 371.504885 5.561485 0.394579
    200000 1 1.055789 531.637407 371.524263 5.567000 0.402265
  • For the input lattice basis in general form, Table II compares the time cost of each algorithm for a fixed size $n=2\times 10^{5}$, as the difference $\Delta=|a_{1}/b_{1}-a_{2}/b_{2}|\approx 10^{-\delta}$ varies, where $n$ denotes the maximum number of decimal digits in $a_{1},a_{2},b_{1}$ and $b_{2}$. The last column, $\kappa$, represents the proportion of differing terms in the continued fraction expansions of $a_{1}/b_{1}$ and $a_{2}/b_{2}$, counted from right to left, relative to the total number of terms in the expansions. Specially, the first $\kappa=1$ indicates that only the last term of the continued fraction expansions of $a_{1}/b_{1}$ and $a_{2}/b_{2}$ differs. Notably, as $\Delta$ decreases, $\delta$ increases and $\kappa$ decreases, indicating that $a_{1}/b_{1}$ and $a_{2}/b_{2}$ become closer to being equal. This further implies that the lattice basis vectors $\mathbf{a}=(a_{1}\ a_{2})^{T}$ and $\mathbf{b}=(b_{1}\ b_{2})^{T}$ approach linear dependence. As shown in Table II, the closer the two input lattice basis vectors are to being linearly dependent ($\delta$ increases), the higher the computational time required for all algorithms. Additionally, the following observations can be made:
    • (i)
[…]
    • (ii)
[…]
    • (iii)
      For our proposed fundamental reduction algorithm $\mathbf{CrossEuc}$ and its optimized variant $\mathbf{HVecSBP}$, the runtime of the latter constitutes approximately $39.5\%$-$19.9\%$ of the former as the linear dependence between input lattice basis vectors decreases – achieving a $2.5$ to $5$-fold efficiency gain. Strikingly, $\mathbf{CrossEuc}$ itself outperforms Yap’s optimized algorithm $\mathbf{HalfGaussianSBP}$, requiring only $14.9\%$-$36.4\%$ of the latter’s runtime under diminishing linear dependence. This advantage is particularly pronounced for smaller-scale lattice bases (characterized by parameter $n$).
TABLE II: Running time of various algorithms for the $\ell_{2}$-shortest basis with varying $\delta$ (in seconds)
$\delta$  CrossEuc   CRS  LagRed  HalfGaussianSBP  HVecSBP  $\kappa$
400000 2.187857 1061.482896 732.066229 14.621288 0.865833 1
350000 2.039634 1029.768712 675.253707 9.911299 0.705745 12.5%
300000 1.963530 953.469648 640.119371 8.779507 0.597472 25%
250000 1.760551 890.452219 595.714993 7.692156 0.530357 37.5%
200000 1.541461 820.747442 564.586658 6.102788 0.430760 50%
150000 1.361687 705.880787 461.356443 4.899072 0.344573 62.5%
100000 1.059967 573.571627 395.397606 3.676309 0.260802 75%
50000 0.753551 416.365192 283.668019 2.415546 0.168508 87.5%
20000 0.571839 306.532068 206.223716 1.718590 0.131812 95%
10000 0.488057 263.113941 178.243286 1.572602 0.103363 97.5%
2000 0.458032 227.192431 159.118865 1.255080 0.090067 99.5%
1000 0.446958 224.975630 158.739315 1.243584 0.087849 99.75%
200 0.439207 223.499143 153.864744 1.234583 0.085077 99.95%
100 0.432651 222.870675 153.261555 1.233634 0.084719 99.975%
20 0.423252 222.065486 152.929364 1.232835 0.084328 99.995%
0 0.422037 221.981643 152.823700 1.231388 0.084183 100%
VI-B2 Experimental result and analysis for the $\ell_{\infty}$-shortest basis
  • For the input lattice basis in HNF, TABLE III compares the time cost of each algorithms as the variance of $n_{2}$, where $n_{1}=10^{6}$ denotes the number of decimal digits in $a,b$, and $n_{2}$ represents the number of decimal digits in $c$. These results highlight HVecSBP’s superior performance in handling HNF-structured bases, particularly in scenarios with highly imbalanced parameter lengths ($n_{2}\ll n_{1}$). Such scenarios are common in cryptographic applications, such as lattice-based attacks, where HVecSBP achieves at least 13.5×13.5$\times$13.5 × efficiency improvement over previous approaches.
    TABLE III: Running time of various algorithms for the $\ell_{\infty}$-shortest basis with varying $n_{2}$ (in seconds)
    length (dec) $n_{1}$  length (dec) $n_{2}$  HVecSBP  CrossEuc/ParEuc  GolEuc
    1000000 1000000 0.006217 0.002445 0.002459
    1000000 800000 0.426119 10.279363 49.977750
    1000000 600000 0.832923 18.195970 90.955213
    1000000 400000 1.162091 26.268203 121.361453
    1000000 200000 1.607177 27.132104 141.404999
    1000000 100000 1.947272 28.337137 146.895788
    1000000 80000 1.919498 28.082500 147.245142
    1000000 60000 1.951394 28.259035 148.577950
    1000000 40000 1.976183 28.483309 148.651564
    1000000 20000 2.039936 28.457997 148.667026
    1000000 10000 2.061487 29.026893 150.630187
    1000000 5000 2.114340 29.114340 150.116029
    1000000 1000 2.153284 29.424312 150.084888
    1000000 100 2.164948 29.353622 150.895833
    1000000 1 2.168530 29.583845 150.976715
  • For the input lattice basis in general form, Table IV compares the time cost of each algorithm for a fixed size $n=10^{6}$, as the difference $\Delta=|a_{1}/b_{1}-a_{2}/b_{2}|\approx 10^{-\delta}$ varies, where $n$ denotes the maximum number of decimal digits in $a_{1},a_{2},b_{1}$ and $b_{2}$. The meaning of the parameter $\kappa$ is same as that in Table II. Notably, as $\Delta$ decreases, $\delta$ increases and $\kappa$ decreases, suggesting that $\frac{a_{1}}{b_{1}}$ and $\frac{a_{2}}{b_{2}}$ become increasingly similar. This, in turn, implies that the lattice basis vectors $\mathbf{a}=(a_{1}\ a_{2})^{T}$ and $\mathbf{b}=(b_{1}\ b_{2})^{T}$ are approaching linear dependence. As shown in Table IV, the following observations can be made:
    • (i)
      For the three fundamental algorithms without HGCD optimization—CrossEuc, EEA-HNF-ParEuc, and GolEuc—the closer the two input lattice basis vectors are to being linearly dependent, the more computational time is required by all three. Meanwhile, our proposed algorithm, CrossEuc, consistently outperforms the other two. Compared to GolEuc, CrossEuc demonstrates a relatively stable efficiency advantage that is largely unaffected by the linear dependency of the input lattice basis vectors. Specifically, the time cost of CrossEuc is approximately 20% of that of GolEuc, representing a 5× speedup. In contrast, when compared to EEA-HNF-ParEuc, the efficiency gain of CrossEuc is highly sensitive to the linear dependency of the input vectors. As the degree of dependency decreases (i.e., as $\delta$ decreases), CrossEuc achieves a speedup ranging from 4×4$\times$4 × to 18×18$\times$18 ×, with the advantage becoming increasingly significant.
    • (ii)
      For the HGCD-optimized algorithms—HGCD-HNF-ParEuc, HGCD-HNF-HVecSBP, and our algorithm HVecSBP, we have (1) When the continued fraction expansions of $a_{1}/b_{1}$ and $a_{2}/b_{2}$ differ only in the last term, HGCD-HNF-ParEuc achieves the highest efficiency, followed by HGCD-HNF-HVecSBP. The time cost of HVecSBP is approximately 2×2$\times$2 × that of the other two algorithms, making it the least efficient. This is intuitive, as the input lattice basis vectors nearly degenerate into two integers, making integer-based HGCD processing more efficient. The vectorized HVec algorithm, which handles two dimensions, naturally incurs double the time cost. (2) As linear dependence between lattice basis vectors diminishes ($\delta$ decreases)—specifically when continued fraction expansions of $a_{1}/b_{1}$ and $a_{2}/b_{2}$ diverge in approximately the last 27.5% of terms—our algorithm $\mathbf{HVecSBP}$ achieves dominant performance. Beyond this threshold, its efficiency advantage scales progressively: accelerating from marginal parity 1×1$\times$1 × to 13-fold speedup over $\mathbf{HGCD\text{-}HNF\text{-}HVecSBP}$, and from 2×2$\times$2 × to 179×179$\times$179 × acceleration against $\mathbf{HGCD\text{-}HNF\text{-}ParEuc}$, with gains intensifying at smaller $\delta$ values.
    • (iii)
      For our proposed fundamental reduction algorithm, $\mathbf{CrossEuc}$, and its optimized variant, $\mathbf{HVecSBP}$, the runtime of $\mathbf{HVecSBP}$ is approximately 4.2% to 8.5% of that of $\mathbf{CrossEuc}$ as the linear dependence between the input lattice basis vectors decreases—resulting in an efficiency gain of roughly $\times1$ to 23×23$\times$23 ×.
TABLE IV: Running time of various algorithms for the $\ell_{\infty}$-shortest basis with varying $\delta$ varies (in seconds)
$\delta$  HVecSBP   HGCD-HNF-HVecSBP  HGCD-HNF-ParEuc  CrossEuc  EEA-HNF-ParEuc  GolEuc  $\kappa$
2000000 4.634715 2.474747 2.335431 54.612207 250.625570 269.501145 1
1750000 3.759720 2.881473 3.509663 53.949515 231.295727 262.973879 12.5%
1500000 3.281029 3.130119 7.128568 53.507613 229.640975 249.979270 25%
1450000 3.177414 3.165364 8.725851 51.548659 210.162490 249.096180 27.5%
1400000 3.092249 3.232044 9.014794 50.038115 208.529073 248.094480 30%
1250000 2.765591 3.551487 14.045340 45.819027 204.894112 231.718764 37.5%
1000000 2.412870 4.063529 21.208691 40.766721 200.351373 206.916571 50%
750000 1.855896 4.274508 31.767607 35.590796 198.790794 173.002045 62.5%
500000 1.417675 4.729311 44.924310 29.244469 197.095717 139.971889 75%
250000 0.917597 5.205109 59.637123 19.667034 196.227950 97.178470 87.5%
100000 0.685130 5.302298 70.196164 14.644298 196.028048 71.587244 95%
50000 0.568257 5.336118 74.210315 13.054622 196.026205 62.898395 97.5%
10000 0.467744 5.331523 75.565236 11.143312 195.498931 54.231681 99.5%
5000 0.459721 5.425184 76.523376 10.757789 195.399899 53.463294 99.75%
1000 0.448851 5.505665 76.635827 10.684234 195.360483 52.133793 99.95%
500 0.447047 5.662536 76.843067 10.612306 195.329952 51.892477 99.975%
100 0.439681 5.689787 77.879209 10.525084 194.966942 51.416951 99.995%
0 0.437364 5.701985 78.296140 10.396514 194.718627 51.037448 100%
VII Conclusion
This paper introduces a newly defined reduced basis for two-dimensional lattices and develops a foundamental reduction algorithm, CrossEuc, which effectively solves the SVP and SBP in two-dimensional lattices. Furthermore, we extend the integer version of the HGCD algorithm and propose a vector-based HVec algorithm, providing detailed implementation and analysis. This leads to an optimized version, HVecSBP, which further accelerates the performance of CrossEuc. Finally, we conduct extensive experiments to evaluate the practical performance of our newly designed algorithms.
Appendix A The optimized Lagrange algorithm under the $\ell_{\infty}$ metric GolEuc.
Algorithm 6 GloEuc[3]
0: A basis $[\mathbf{a}\ \mathbf{b}]$.
0: An $\ell_{\infty}$-shortest basis for $\mathcal{L}([\mathbf{a}\ \mathbf{b}])$
1:If $\|\mathbf{a}\|>\|\mathbf{b}\|,$ then swap ($\mathbf{a},\mathbf{b}$)
2:If $\|\mathbf{a}-\mathbf{b}\|>\|\mathbf{a}+\mathbf{b}\|,$ then $\mathbf{b}:=-\mathbf{b}$
3:If $\|\mathbf{b}\|\leq\|\mathbf{a}-\mathbf{b}\|,$ then return $[\mathbf{a}\ \mathbf{b}]$
4:If $\|\mathbf{a}\|\leq\|\mathbf{a}-\mathbf{b}\|,$ then goto loop
5:If $\|\mathbf{a}\|=\|\mathbf{b}\|,$ then return $[\mathbf{a}\ \mathbf{a}-\mathbf{b}]$
6:$[\mathbf{a}\ \mathbf{b}]:=[\mathbf{b}-\mathbf{a}\ \mathbf{a}]$
7:loop
8:  If $\frac{b_{1}}{a_{1}},\frac{b_{2}}{a_{2}}\neq 0$ and $sgn\left(\frac{b_{1}}{a_{1}}\right)=sgn\left(\frac{b_{2}}{a_{2}}\right)$ then
9:   Find $\mu\in\left\{\left\lceil\frac{|b_{1}|+|b_{2}|}{|a_{1}|+|a_{2}|}\right\rceil,\left\lfloor\frac{|b_{1}|+|b_{2}|}{|a_{1}|+|a_{2}|}\right\rfloor\right\},$ s.t. $\|\mathbf{b}-\mu\mathbf{a}\|$ is minimal
10:  Else
11:   Find $\mu\in\left\{\left\lceil\frac{\big{|}|b_{2}|-|b_{1}|\big{|}}{|a_{2}|+|a_{1}|}\right\rceil,\left\lfloor\frac{\big{|}|b_{2}|-|b_{1}|\big{|}}{|a_{2}|+|a_{1}|}\right\rfloor\right\},$ s.t. $\|\mathbf{b}-\mu\mathbf{a}\|$ is minimal
12:  $\mathbf{b}:=\mathbf{b}-\mu\mathbf{a}$
13:  If $\|\mathbf{a}-\mathbf{b}\|>\|\mathbf{a}+\mathbf{b}\|$ then $\mathbf{b}:=-\mathbf{b}$
14:  Swap($\mathbf{a},\mathbf{b}$)
15:  If $[\mathbf{a}\ \mathbf{b}]$ is Lagrange-reduced, then return $[\mathbf{a}\ \mathbf{b}]$
16:  goto loop
Appendix B Lagrange’s reduction algorithm under the $\ell_{2}$ metric LagRed
Algorithm 7 LagRed [9]
0: A basis $[\mathbf{a}\ \mathbf{b}]$.
0: An $\ell_{2}$-shortest basis for $\mathcal{L}([\mathbf{a}\ \mathbf{b}])$
1:If $\|\mathbf{a}\|_{2}>\|\mathbf{b}\|_{2},$ then swap ($\mathbf{a},\mathbf{b}$)
2:If $\|\mathbf{a}-\mathbf{b}\|_{2}>\|\mathbf{a}+\mathbf{b}\|_{2},$ then $\mathbf{b}:=-\mathbf{b}$
3:If $\|\mathbf{b}\|_{2}\leq\|\mathbf{a}-\mathbf{b}\|_{2},$ then return $[\mathbf{a}\ \mathbf{b}]$
4:If $\|\mathbf{a}\|_{2}\leq\|\mathbf{a}-\mathbf{b}\|_{2},$ then goto loop
5:If $\|\mathbf{a}\|_{2}=\|\mathbf{b}\|_{2},$ then return $[\mathbf{a}\ \mathbf{a}-\mathbf{b}]$
6:$[\mathbf{a}\ \mathbf{b}]:=[\mathbf{b}-\mathbf{a}\ \mathbf{a}]$
7:loop
8:  $\mu:=\left\lceil\frac{<\mathbf{a},\mathbf{b}>}{||\mathbf{a}||_{2}^{2}}\right\rfloor$
9:  $\mathbf{b}:=\mathbf{b}-\mu\mathbf{a}$
10:  If $\|\mathbf{a}-\mathbf{b}\|_{2}>\|\mathbf{a}+\mathbf{b}\|_{2}$ then $\mathbf{b}:=-\mathbf{b}$
11:  Swap($\mathbf{a},\mathbf{b}$)
12:  If $[\mathbf{a}\ \mathbf{b}]$ is Lagrange-reduced, then return $[\mathbf{a}\ \mathbf{b}]$
13:  goto loop
Appendix C Yap’s coherent remainder sequence Reduction Algorithm CRS
Algorithm 8 CRS [15]
0: An admissible basis $[\mathbf{a}\ \mathbf{b}]$ refers to a pair of vectors $\mathbf{a}$ and $\mathbf{b}$ that form an acute angle and satisfy $\|\mathbf{a}\|_{2}\geq\|\mathbf{b}\|_{2}$.
0: An $\ell_{2}$-shortest basis for $\mathcal{L}([\mathbf{a}\ \mathbf{b}])$
1:While(true)
2:  $\mu:=\left\lfloor\frac{<\mathbf{a},\mathbf{b}>}{||\mathbf{b}||_{2}^{2}}\right\rfloor$
3:  If $||\mathbf{a}-\mu\mathbf{b}||_{2}=||\mathbf{a}||_{2}$, break
4:  $[\mathbf{a}\ \mathbf{b}]:=[\mathbf{b}\ \mathbf{a}-\mu\mathbf{b}]$
5:If $||\mathbf{a}-\mathbf{b}||_{2}<||\mathbf{b}||_{2}$, then $[\mathbf{a}\ \mathbf{b}]:=[\mathbf{b}\ \mathbf{a}-\mathbf{b}]$
[…]
7:Return $[\mathbf{a}\ \mathbf{b}]$
Appendix D Yap’s Half-Gaussian Algorithm Half-Gaussian
Algorithm 9 Half-Gaussian [15]
0: An admissible basis $[\mathbf{a}\ \mathbf{b}]$ refers to a pair of vectors $\mathbf{a}$ and $\mathbf{b}$ that form an acute angle and satisfy $\|\mathbf{a}\|_{2}\geq\|\mathbf{b}\|_{2}$.
0: Regular matrix $\mathbf{M^{*}}$ that reduce $[\mathbf{a}\ \mathbf{b}]$ to $[\mathbf{a}^{*}\ \mathbf{b}^{*}]$ such that $[\mathbf{a}^{*}\ \mathbf{b}^{*}]$ either satisfy $(\log\|\mathbf{a^{*}}\|_{2}\geq(\log\|\mathbf{a}\|_{2}/2)+c>\log\|\mathbf{b^{*}}\|_{2}$ (also called straddles) or is terminal with $\log\|\mathbf{b^{*}}\|_{2}\geq\log\|\mathbf{a}\|_{2}+c$, $c$ is a very large positive integer.
1:If $\cos(\mathbf{a},\mathbf{b})\geq 1/2$, return $\mathbf{M^{*}}:=\mathbf{E}$
$n:=\left\lceil\log||\mathbf{a}||_{2}\right\rceil,m:=\left\lceil n/2\right\rceil$
3:If $\log||\mathbf{b}||_{2}<n-(m/2)+c$ or $m\leq 8$, then $\mathbf{M}_{1}^{*}:=\mathbf{E}$ and goto step $20$
4:$(\mathbf{a_{0}},\mathbf{b_{0}}):=(\left\lceil\mathbf{a}/2^{m}\right\rceil,\left\lfloor\mathbf{b}/2^{m}\right\rfloor)$
5:If $||\mathbf{a_{0}}||_{2}=||\mathbf{b_{0}}||_{2}$, then $\mathbf{M}_{1}^{*}:=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ and goto step $20$
6:$\mathbf{M_{1}}:=\textbf{Half-Gaussian}(\mathbf{a_{0}},\mathbf{b_{0}})$
7:$(\mathbf{a^{\prime}},\mathbf{b^{\prime}}):=(\mathbf{a},\mathbf{b})\mathbf{M_{1}^{-1}}$
8:$(\mathbf{a_{0}^{\prime}},\mathbf{b_{0}^{\prime}}):=(\mathbf{a_{0}},\mathbf{b_{0}})\mathbf{M_{1}^{-1}}$
9:If $\log||\mathbf{a}_{0}^{\prime}||_{2}\geq\log||\mathbf{a}_{0}||_{2}-(m/2)+c>\log||\mathbf{b}_{0}^{\prime}||_{2}$
10:  If $\log||\mathbf{a}^{\prime}||_{2}\leq\log||\mathbf{b}^{\prime}||_{2}$, then $\mathbf{M}_{1}^{*}:=\begin{pmatrix}0&1\\ 1&-q_{k}\end{pmatrix}\mathbf{M_{1}}$
11:  If $\cos(\mathbf{a^{\prime}},\mathbf{b^{\prime}})<0$
12:   $\mathbf{a^{\prime\prime}}:=\mathbf{a^{\prime}}+\mathbf{b^{\prime}}$
13:   If $q_{k}>1$
14:    If $\log||\mathbf{a^{\prime}}||_{2}>\log||\mathbf{a^{\prime\prime}}||_{2}$, let $q_{k}:=q_{k}-1$ and update $\mathbf{M_{1}}$ to $\mathbf{M}_{1}^{*}$
15:    Else let $q_{k}:=q_{k}-1$ and update $\mathbf{M_{1}}$ to $\mathbf{M}_{1}^{*}$, then make one toggle: $\mathbf{M}_{1}^{*}:=\begin{pmatrix}1&1\\ 1&0\end{pmatrix}\mathbf{M}_{1}^{*}$
16:   Else
17:    If $\log||\mathbf{a^{\prime\prime}}||_{2}>\log||\mathbf{a^{\prime}}||_{2}$, backup one step and update $\mathbf{M_{1}}$ to $\mathbf{M}_{1}^{*}$
18:    Else backup two steps and update $\mathbf{M_{1}}$ to $\mathbf{M}_{1}^{*}$
19:Else $\mathbf{M}_{1}^{*}:=\mathbf{M_{1}}$
20:$(\mathbf{a_{1}^{*}},\mathbf{b}_{1}^{*}):=(\mathbf{a},\mathbf{b})\mathbf{M_{1}^{*}}^{-1}$
[…]
22:Else do one step CRS-red and compute the elementary matrix $\mathbf{M}$ to transform $(\mathbf{a}_{1}^{*},\mathbf{b}_{1}^{*})$ to $(\mathbf{b}_{1}^{*},\mathbf{c}_{1}^{*})$, let $(\mathbf{x},\mathbf{y}):=(\mathbf{b}_{1}^{*},\mathbf{c}_{1}^{*})$
23:If $(\mathbf{x},\mathbf{y})$ straddles $(n/2)+c$, return $\mathbf{M}\mathbf{M}_{1}^{*}$
24:$n^{\prime}:=\left\lceil\log||\mathbf{x}||_{2}\right\rceil,m^{\prime}:=2(n^{\prime}-m)$
25:If $\log||\mathbf{y}||_{2}<n^{\prime}-(m^{\prime}/2)+c$ or $m^{\prime}\leq 8$, then $\mathbf{M_{2}^{*}}:=\mathbf{E}$ and goto step $42$
26:$(\mathbf{x_{0}},\mathbf{y_{0}}):=(\left\lceil\mathbf{x}/2^{m^{\prime}}\right\rceil,\left\lfloor\mathbf{y}/2^{m^{\prime}}\right\rfloor)$
27:If $||\mathbf{x_{0}}||_{2}=||\mathbf{y_{0}}||_{2}$, then $\mathbf{M_{2}^{*}}:=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ and goto step $42$
28:$\mathbf{M_{2}}:=\textbf{Half-Gaussian}(\mathbf{x_{0}},\mathbf{y_{0}})$
29:$(\mathbf{x^{\prime}},\mathbf{y^{\prime}}):=(\mathbf{x},\mathbf{y})\mathbf{M_{2}^{-1}}$
30:$(\mathbf{x_{0}^{\prime}},\mathbf{y_{0}^{\prime}}):=(\mathbf{x_{0}},\mathbf{y_{0}})\mathbf{M_{2}^{-1}}$
31:If $\log||\mathbf{x}_{0}^{\prime}||_{2}\geq\log||\mathbf{x_{0}}||_{2}-(m^{\prime}/2)+c>\log||\mathbf{y}_{0}^{\prime}||_{2}$
32:  If $\log||\mathbf{x}^{\prime}||_{2}\leq\log||\mathbf{y}^{\prime}||_{2}$, then $\mathbf{M_{2}^{*}}:=\begin{pmatrix}0&1\\ 1&-q_{k}\end{pmatrix}\mathbf{M}_{2}$
33:  If $\cos(\mathbf{x^{\prime}},\mathbf{y^{\prime}})<0$
34:   $\mathbf{x^{\prime\prime}}:=\mathbf{x^{\prime}}+\mathbf{y^{\prime}}$
35:   If $q_{k}>1$
36:    If $\log||\mathbf{x^{\prime}}||_{2}>\log||\mathbf{x^{\prime\prime}}||_{2}$, let $q_{k}:=q_{k}-1$ and update $\mathbf{M_{2}}$ to $\mathbf{M_{2}^{*}}$
37:    Else let $q_{k}:=q_{k}-1$ and update $\mathbf{M_{2}}$ to $\mathbf{M_{2}^{*}}$, then make one toggle: $\mathbf{M_{2}^{*}}:=\begin{pmatrix}1&1\\ 1&0\end{pmatrix}\mathbf{M_{2}^{*}}$
38:   Else
39:    If $\log||\mathbf{x^{\prime\prime}}||_{2}>\log||\mathbf{x^{\prime}}||_{2}$, backup one step and update $\mathbf{M_{2}}$ to $\mathbf{M_{2}^{*}}$
40:    Else backup two steps and update $\mathbf{M_{2}}$ to $\mathbf{M_{2}^{*}}$
41:Else $\mathbf{M_{2}^{*}}:=\mathbf{M_{2}}$
42:return $\mathbf{M^{*}}:=\mathbf{M_{2}^{*}}\mathbf{M}\mathbf{M}_{1}^{*}$
Algorithm 10 HalfGaussianSBP($\mathbf{a},\mathbf{b}$)
0:  A base $[\mathbf{a}\ \mathbf{b}]$ with $\mathbf{a}=(a_{1}\ a_{2})^{T}$, $\mathbf{b}=(b_{1}\ b_{2})^{T}$
0:  An $\ell_{2}$-shortest basis $[\mathbf{a}\ \mathbf{b}]$
1:  $\mathbf{M}:=\mathbf{I}$
2:  While (true)
3:   If $||\mathbf{a}||_{2}<||\mathbf{b}||_{2}$, Swap($\mathbf{a},\mathbf{b}$)
4:   If $\cos(\mathbf{a},\mathbf{b})<0$, $\mathbf{b}:=-\mathbf{b}$
5:   $\mathbf{M}:=\textbf{Half-Gaussian}(\mathbf{a},\mathbf{b})$
6:   $(\mathbf{a},\mathbf{b}):=(\mathbf{a},\mathbf{b})\mathbf{M^{-1}}$
7:   $\mu:=\left\lfloor\frac{<\mathbf{a},\mathbf{b}>}{||\mathbf{b}||_{2}^{2}}\right\rfloor$
8:   If $||\mathbf{a}-\mu\mathbf{b}||_{2}=||\mathbf{a}||_{2}$, break
9:   $[\mathbf{a}\ \mathbf{b}]:=[\mathbf{b}\ \mathbf{a}-\mu\mathbf{b}]$
10:  If $||\mathbf{a}-\mathbf{b}||_{2}<||\mathbf{b}||_{2}$, then $[\mathbf{a}\ \mathbf{b}]:=[\mathbf{b}\ \mathbf{a}-\mathbf{b}]$
11:  $[\mathbf{a}\ \mathbf{b}]:=[\mathbf{b}\ \mathbf{a}-\left\lceil\frac{<\mathbf{a},\mathbf{b}>}{||\mathbf{b}||_{2}^{2}}\right\rfloor\mathbf{b}]$
12:  Return $[\mathbf{a}\ \mathbf{b}]$
Lihao Zhao received the B.E. degree in information security from Qingdao University in 2023. He is currently pursuing the M.S. degree in the School of Computer Science and Technology of Qingdao University. His research interests include lattice algorithms and privacy computing.
Chengliang Tian received the B.S. and M.S. degrees in mathematics from Northwest University, Xi’an, China, in 2006 and 2009, respectively, and the Ph.D. degree in information security from Shandong University, Ji’nan, China, in 2013. He held a post-doctoral position with the State Key Laboratory of Information Security, Institute of Information Engineering, Chinese Academy of Sciences, Beijing. He is currently with the College of Computer Science and Technology, Qingdao University, as an Associate Professor. His research interests include lattice-based cryptography and cloud computing security.
Jingguo Bi received the B.Sc. and Ph.D. degrees in information security from Shandong University, Jinan, China, in 2007 and 2012, respectively.He is currently an Associate Researcher with the School of Cyberspace Security, Beijing University of Posts and Telecommunications, Beijing, China. His research interests are public key cryptography, cloud computing, and post-quantum cryptography.
Guangwu Xu (Senior Member) received his Ph.D. in mathematics from SUNY Buffalo. He is now with the School of Cyber Science and Technology, Shandong University, China. His research interests include cryptography, arithmetic number theory, compressed sensing, algorithms, and functional analysis.
Jia Yu (Member, IEEE) received the B.S. and M.S. degrees from the School of Computer Science and Technology, Shandong University, Jinan, China, in 2003 and 2000, respectively, and the Ph.D. degree from the Institute of Network Security, Shandong University in 2006. He is a Professor with the College of Computer Science and Technology, Qingdao University, Qingdao, China. He was a Visiting Professor with the Department of Computer Science and Engineering, State University of New York at Buffalo, Buffalo, NY, USA, from November 2013 to November 2014. His research interests include cloud computing security, key evolving cryptography, digital signature, and network security